### v2

# Detecting Hidden Triggers: Mapping Non-Markov Reward Functions to Markov

Abstract
Many Reinforcement Learning algorithms assume a Markov reward function to guarantee optimality. However, not all reward functions are known to be Markov. This paper proposes a framework for mapping non-Markov reward functions into equivalent Markov ones by learning a Reward Machine - a specialized reward automaton. Unlike the general practice of learning Reward Machines, we do not require a set of high-level propositional symbols from which to learn. Rather, we learn hidden triggers directly from data that encode them. Given their ability to model reward dependencies, we demonstrate the importance of learning Reward Machines versus their Deterministic Finite-State Automata counterparts. We formalize this distinction in our learning objective. Our mapping process is constructed as an Integer Linear Programming problem. We prove that our mappings provide consistent expectations for the underlying process. We empirically validate our approach by learning black-box non-Markov Reward functions in the Officeworld Domain. Additionally, we demonstrate the effectiveness of learning dependencies between rewards in a new domain, Breakfastworld.
1 Introduction
Reinforcement Learning (RL) is traditionally formulated as a sequential decision-making task modeled under a Markov Decision Process (MDP) [4]. This implies that both the reward and transition functions are Markov, depending solely on the current state and action. However, in complex environments, these functions are not always known and cannot always be assumed to be Markov. In fact, it has been shown that there are behaviors that cannot be encoded by any Markov reward function under certain state representations [1]. To address non-Markov dynamics, researchers often extend the state representation to enforce the Markov property, though this is usually at the cost of overcomplicating the state representation or making it excessively sparse. For example, Mnih et al. [19] stacked subsequent frames from the Atari games [3] to encode a suitable representation for training Deep Q-Networks. Similarly, deep, recurrent models have been used encode state histories for similar purposes [15, 11]. However, these sparse, deep models obscure the interpretation of rewards. To better understand them it is crucial to examine their dynamics under less sparse and more interpretable forms.
A common strategy to represent non-Markov rewards in an interpretable format is to use the finite states of an automata to represent hidden or historical features that render the representation Markov. Historically, this has been accomplished with Deterministic Finite-State Automata (DFA) [5, 13, 12], but more recently, specialized reward automata, Reward Machines (RMs), have been used [32, 22, 23, 10]. We highlight two important distinctions between DFAs and RMs: 1.) DFAs are inherently less expressive than RMs as they only accept or reject traces. Consequently, when rewards are complex, multiple DFAs must be learned [13]. In contrast, RMs represent complex reward behavior in a single automaton. 2.) DFAs are only capable of representing patterns over $H=(S\times A)^{*}$, the history of states and actions. Conversely, RMs are capable of representing patterns over $H=(S\times A\times\mathbb{R})^{*}$, the history of states, actions, and rewards. This affords us a unique learning objective where rewards serve as memory allowing us to model reward dependencies.
However, prior work for learning RMs assumes access to $P$, the set of high-level propositional symbols the RM operates over, and $L:S\times A\times S\rightarrow 2^{P}$, a labeling function mapping transitions to instantiations in $P$. This makes much of this prior work ungeneralizable. We instead infer $L$ by identifying minimal patterns in $H=(S\times A\times\mathbb{R})^{*}$, called hidden triggers, that differentiate reward outcomes. Rather than learn the RM directly, we map observed non-Markov rewards onto the cross-product of $U$, the state space of a candidate RM, and $S$, the observed state space, to form a Markov representation called the Abstract Reward Markov Decision Process (ARMDP). We infer our ARMDPs, and thus its decomposable RM, by solving an Integer Linear Program (ILP), lending itself to powerful off-the-shelf discrete optimization solvers.
Contributions
We introduce a novel algorithm for mapping non-Markov reward functions to equivalent Markov ones. We show how leveraging $H=(S\times A\times\mathbb{R})^{*}$ in our learning objective can dramatically expedite learning in cases with multiple interdependent reward signals. We validate our approach in several ways. First, we demonstrate our ability to map non-Markov rewards sampled from black-box RMs in the Officeworld domain, originally introduced in the first RM paper [16]. Secondly, we evaluate the representative power of the ARMDP by extending DQN agents with ARMDP features to compare their learning profiles against various sparse Recurrent DQN (RDQN) variants. Finally, we validate our approach to represent complex reward function behavior in a new domain, Breakfastworld, featuring black-box RMs with interdependent reward signals. Importantly, we show how by leveraging $H=(S\times A\times\mathbb{R})^{*}$ we can expedite learning in these cases. Our contributions are three-fold:
  1. 1.
    We introduce a novel algorithm for RM learning without assumed access to $P$ and $L$,
  2. 2.
    We prove that the ARMDP is a suitable representation for maximizing reward expectations under non-Markov rewards,
  3. 3.
    We demonstrate the effectiveness of learning RMs under multiple interdependent reward signals.
2 Related Work
Decision-making with non-Markov reward functions is often represented under the Non-Markov Reward Decision Process (NMRDP) [27]. An NMRDP is similar to an MDP except that the reward function, $R:H\rightarrow\mathbb{R}$, maps state and action histories, $H=(S\times A)^{*}$, to reward values. However, as $H$ can become unwieldy in large state and action spaces, more succinct representations have been encoded using automata. The intuition of this approach lies in the fact that automata are finitely represented and offer a level of interpretation due to their structured form and symbolic representation.
Traditionally, the automata of choice has been the DFA, yielding promising results in identifying Markov representations for NMRDPs [5, 13, 12]. However, more recently, specialized reward automata, RMs, have been used to represent NMRDP behavior [16]. While DFAs operate over an input alphabet and emit true or false over traces, RMs extend this expressiveness by emitting reward values along each transition. RMs are particularly appealing as they offer succinct encoding of non-Markov reward behavior and task decomposition. They have also been extended by a broad range of RL algorithms that exploit RM structure to learn effective or optimal (in the discrete case) policies [7]. These include but are not limited to, specialized Q-Learning (QRM), Counterfactual experiences (CRM), Hierarchical Reinforcement Learning (HRM) and Reward Shaping (RSRM) for RMs [7, 17].
Whether using a DFA or an RM, various strategies for learning them exist. One prominent strategy for learning automata is to use Angluin’s L* algorithm [2]. L* has been applied for both DFA [13] and RM [22, 23, 10] learning. Other strategies include Inductive Logic Answer Set Programming (ILASP) [12] for DFAs and SAT solvers [32] or Tabu search [29] for RMs. Non-automata-based learning of non-Markov rewards has been accomplished via Linear Temporal Logic ($LTL$), [18, 5, 6]. LTL can also be later translated into automata. Camacho et al. provide a helpful pipeline for transforming LTL formulae into RMs [7].
A notable limitation of these prior works is that they assume access to a set of high-level symbols (the input language for the automata) and the relationship of the underlying state space to those symbols. In RM literature these are denoted $P$ and $L:S\times A\times S\rightarrow 2^{P}$ where $P$ is a set of symbols, separate from the NMRDP, and $L$ is a mapping of transitions from the NMRDP to instantiations in $P$, respectively. A similar assumption is often made for DFA learning, with exceptions being the work of Gaon and Brafman [13] as well as Christoffersen et al. [9] who learn symbolic representations from $H:(S\times A)^{*}$ to build their automata. Gaon and Brafman learn their DFA using L*, however, their approach was shown to be sample inefficient. Christoffersen et al. improved upon this by solving a discrete optimization problem for DFA learning, as outlined by Shvo et al. [24], which regularizes automata based on their size to better handle noise and enhance generalization. We improve upon these approaches by, instead, learning an RM without assumed access to $P$ and $L$. We show how by learning RMs we can model reward dependencies in a single automata by representing patterns in $H:(S\times A\times\mathbb{R})^{*}$.
3 Preliminaries
3.1 RL and Non-Markov Rewards
RL is a type of self-supervised learning in which an agent learns to interact with its environment to maximize its long-term reward expectation [25]. It is an iterative process in which the agent receives feedback in the form of rewards or punishments to adjust its behavior. RL is often represented as a sequential decision-making problem formalized using the MDP. An MDP is a 6-tuple $(S,A,T,R,\gamma,\rho)$, where $S$ is the state space, $A$ is the action space, $T:S\times A\times S\rightarrow[0,1]$ is a Markov probabilistic transition function, $R:S\times A\times S\rightarrow\mathbb{R}$ is a Markov reward function mapping states, actions and states to rewards, $\gamma\in[0,1)$ is the discount factor and $\rho:S\rightarrow[0,1]$ is the initial state distribution. We note that NMRDPs are similar to MDPs in that they share $(S,A,T,R,\rho)$, but $R:H\rightarrow\mathbb{R}$ in NMRDPs.
The solution for an MDP is a policy, $\pi:S\rightarrow A$, that maps states to actions. The value of a policy for any state, $s\in S$, at time point $t$ is the expected discounted return defined as follows:
$V_{\pi}(s)=\mathbb{E}\left[\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}\mid s_{t}=s\right]$ (1)
The optimal policy denoted $\pi_{*}$, is determined by solving for the optimal state value function:
$V_{*}(s)=\max_{\pi}V^{\pi}(s)$ (2)
for all $s\in S$. We can write the optimal state-action value function in terms of the state value function as follows:
$Q_{*}(s,a)=\mathbb{E}\left[r_{t+1}+\gamma V_{*}(s_{t+1})\mid s_{t}=s,a_{t}=a\right]$ (3)
for all $s\in S$ and $a\in A$. Hence:
$V_{*}(s)=\max_{a}Q_{*}(s,a)$ (4)
Equation (4) is often solved through a process called Value Iteration [25]. The optimal policy is realized by selecting actions according to $\max_{a}Q_{*}(s,a)$, otherwise known as a greedy policy. A popular model-free approach for learning Equation (4) is Q-Learning [30]:
[…]
where $q(s_{t},a_{t})$ is updated slowly according to $\alpha\in[0,1)$. When updating Equation (5) in an online fashion, typically an $\epsilon$-greedy strategy is employed where the agent follows a greedy policy but chooses a random action $\epsilon-$percent of the time.
Both strategies are guaranteed to converge to the optimal policy in the limit as each state-action pair is visited infinitely often, but only under the assumption that $T$ and $R$ are Markov. Otherwise, expectations have no stable solution.
3.2 Reward Machines
RMs are a specialized type of automata imposed over an underlying state and action space, $S$ and $A$ [16]. RMs operate over a set of propositional symbols, $P$, representing high-level events, separate from $S$. RMs are defined as a 5-tuple, $(U,u_{1},F,\delta_{u},\delta_{r})$, where $U$ is a finite state space, $u_{1}\in U$ is the initial state, $F$ is a set of terminal states where $F\cap U=\emptyset$, $\delta_{u}:U\times 2^{P}\rightarrow U\cup F$ is a transition function and $\delta_{r}:U\times 2^{P}\times U\cup F\rightarrow\mathbb{R}$ is a reward function. A labeling function, $L:S\times A\times S\rightarrow 2^{P}$, connects $S$ and $A$ to the RM by mapping transitions to instantiations in $P$.
Figure 1: The Officeworld domain (a) with four Reward Machines, (b) deliver coffee to office, (c) deliver mail to office, (d) deliver coffee and mail to office and (e) patrol task sequencing A, B, C and D.
RMs are capable of modeling reward dependencies based on histories that can be represented as a regular language (e.g., loops, conditions, and interleaving). RMs were initially demonstrated on the Officeworld domain [16] shown in Figure 1 (a). Tasks (b-e) represent RMs sensitive to $P=\{c,m,o,*,A,B,C,D\}$ where $c$ is a coffee tile, $m$, the mail tile, $o$, the office tile, *, a plant tile, and $A,B,C,D$ are room IDs. $\neg$, $\lor$ and $\land$ are logical operators for negation, or, and and, respectively. The cross-product of $U$ and $S$ is Markov, which, in the discrete case, means we can solve these problems optimally.
4 Approach
We wish to map non-Markov reward signals into succinct Markov representations. We do so by inferring minimal RMs that represent hidden or historical features that render the representation Markov. To make our approach generalizable to practical decision-making scenarios, we drop the assumed access to $P$ and $L$. For simplicity of notation, we drop notation on terminal states, $F$.
Problem Statement
From a set of observed, finite-length trajectories, $T_{o}$, sampled from an NMRDP, $(S,A,T,R)$, we learn a minimal RM $(U,u_{1},\delta_{u},\delta_{r})$ that renders rewards Markov to $S$, $A$ and $U$.
While $R$ in the NMRDP is unknown, we assume it can be sampled from and is representable as an RM. Thus, hereafter, we will refer to $R$ as a black-box RM. Our approach to solving this problem is first to learn an ARMDP, which is simply the MDP cross-product between an RM and an NMRDP, that enables the extraction of the associated RM. Learning of the ARMDP is formulated as an ILP. A minimal RM is an RM with the smallest hidden state space $|U|$, and the least number of transition edges between states.
The following sections are organized as follows. We begin by defining the ARMDP as the cross-product space between an RM and an NMRDP in Section 4.1. Section 4.2 describes a passive learning approach for learning ARMDPs from $T_{o}$, by solving an ILP. We later make this an active learning approach in Section 4.5 by incorporating a Q-Learning agent. Section 4.3 briefly outlines how to extract an RM from an ARMDP. Finally, Section 4.4 describes the theoretical insights of ARMDPs as a proxy for the NMRDP.
4.1 Abstract Reward Markov Decision Process
An ARMDP is an MDP with additional properties sufficient for extracting/learning a valid RM.
Definition 1.
Given an NMRDP $(S,A,T,R)$, RM $(U,u_{1},\delta_{u},\delta_{r})$, and labeler, $L:S\times A\times S\rightarrow 2^{P}$, an ARMDP is an MDP $(\tilde{S},A,\tilde{T},\tilde{R},\gamma,\tilde{\rho})$ satisfying:
  1. 1.
    $\tilde{S}=S\times U$
  2. 2.
    $\{u\mid(s,u)\in\tilde{S},\tilde{\rho}((s,u))>0\}=\{u_{1}\}$
  3. 3.
    $\tilde{T}((s,u),a,(s^{\prime},u^{\prime}))=\begin{cases}T(s,a,s^{\prime})&\text{if\ }\delta_{u}(u,L(s,a,s^{\prime}))=u^{\prime}\\ 0&\text{otherwise}\end{cases}$
  4. 4.
    $\tilde{R}((s,u),a,(s^{\prime},u^{\prime}))=\delta_{r}(u,L(s,a,s^{\prime}),u^{\prime})$.
Property 1 defines the state space of the ARMDP as the cross-product of $S$ (what we observe) and $U$ (what we do not observe). Property 2 imposes an initial state distribution compliant with $u_{1}\in U$ being the initial RM state. Property 3 defines the transition function of the ARMDP as a composition function of $T$ and $\delta_{u}$. Finally, Property 4 defines the reward function of the ARMDP as $\delta_{r}$.
The intuition behind the ARMDP is that if we assume rewards in $T_{o}$ are Markov to $S$ and $A$, conflicts are bound to arise. Consider $\delta_{r}:U\times 2^{P}\times U\rightarrow\mathbb{R}$, for some black-box RM, being rewritten as $\delta_{r}:U\times S\times A\times S\times U\rightarrow\mathbb{R}$ by invoking $L$. Because we do not observe $U$, we lack the specificity to predict $\delta_{r}$. Our goal is to find some candidate ARMDP, that renders rewards Markov to $S$, $A$, and $U$, by identifying hidden triggers in $H=(S\times A\times\mathbb{R})^{*}$ that resolve all reward conflicts in $T_{o}$.
4.2 Formulation
We infer candidate ARMDPs by solving an ILP, that resolves reward conflicts over $T_{o}$. Let, $\tau_{m}\in T_{o}$, be the m-th trace in $T_{o}$ of length $l_{m}$. Further, let $\tau_{m,n}=(s_{n}^{m},a_{n}^{m},r_{n+1}^{m},s_{n+1}^{m})$ be the n-th tuple of $\tau_{m}$. We define a reward conflict to be any two $\tau_{m,n}$ and $\tau_{\tilde{m},\tilde{n}}$ where $(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=(s_{\tilde{n}}^{\tilde{m}},a_{\tilde{n}}^{\tilde{m}},s_{\tilde{n}+1}^{\tilde{m}})$, but $r_{n+1}^{m}\neq r_{\tilde{n}+1}^{\tilde{m}}$.111 While we do not consider stochastic or noisy rewards in this work, we posit reward conflicts might be detected by sampling histories, in these cases. For example, if we define the history of $\tau_{m,n}$ to be $h(\tau_{m,n})=\{(s_{1}^{m},a_{1}^{m},r_{2}^{m},s_{2}^{m}),(s_{2}^{m},a_{2}^{m},r_{3}^{m},s_{3}^{m}),...,(s_{n-1}^{m},a_{n-1}^{m},r_{n}^{m},s_{n}^{m})\}$, then if $\tau_{m,n}$ is observed to conflict with $\tau_{\tilde{m},\tilde{n}}$ we might continually sample over $h(\tau_{m,n})$ and $h(\tau_{\tilde{m},\tilde{n}})$ to assess if their reward distributions are different in some statistical sense (e.g., difference of means). The motivation of our ILP is to map $T_{o}$ onto a valid ARMDP such that we resolve all reward conflicts observed in $T_{o}$. Intuitively, $\tilde{S}$, provides a degree of freedom over $U$, to do just that.
We define the matrix, $O=M\times N\times K\times K$, where $M=|T_{o}|$, $N$ reflects the largest $l_{m}$, $\forall\tau_{m}\in T_{o}$, and $K=|U|$ reflects the size of the finite state space for some candidate RM. Each $O_{m,n}$, then, is a $K\times K$ matrix that lays the foundation for Property 1 in the ARMDP:
$O_{m,n}=\begin{pmatrix}O_{m,n,1,1}&O_{m,n,1,2}&\cdots&O_{m,n,1,K}\\ O_{m,n,2,1}&O_{m,n,2,2}&\cdots&O_{m,n,2,K}\\ \vdots&\vdots&\ddots&\vdots\\ O_{m,n,K,1}&O_{m,n,K,2}&\cdots&O_{m,n,K,K}\end{pmatrix}$
Each $O_{m,n,i,j}\in O_{m,n}$ is a binary variable reflecting the decision to map $s_{n}^{m}$ to RM state $u_{i}\in U$ and $s_{n+1}^{m}$ to RM state $u_{j}\in U$. That is, if $O_{m,n,i,j}=1$, then $\tau_{m,n}=(s_{n}^{m},a_{n}^{m},r_{n+1}^{m},s_{n+1}^{m})$ is mapped onto the ARMDP tuple, $((s_{n}^{m},u_{i}),a_{n}^{m},r_{n+1}^{m}(s_{n+1}^{m},u_{j}))$. We constrain each $O_{m,n}$ to choose a single mapping by enforcing that $\forall m=1,2,...,M,\forall n=1,2,...,l_{m}$:
$\sum_{i=1}^{K}\sum_{j=1}^{K}O_{m,n,i,j}=1$ (6)
We further constrain $O_{m,n}$ in instances where $n=1$ to satisfy the initial state requirement of Property 2 in the ARMDP. That is, $\forall m=1,2,...,M$:
$\sum_{j=1}^{K}O_{m,1,1,j}=1$ (7)
Next, we ensure the continuity of each trajectory in the mapping process. That is, we encode agreement between subsequent mappings of $O_{m,n}$ and $O_{m,n+1}$. To this end, we enforce the following constraint $\forall m=1,2,...,M,\forall n=1,2,...,l_{m}-1,\forall j=1,2,...,K$:
$\sum_{i=1}^{K}O_{m,n,i,j}=\sum_{i=1}^{K}O_{m,n+1,j,i}$ (8)
Note that we swap indices $j$ and $i$ on the right-hand side. Equation (8) simply forces $s_{n+1}^{m}$ to be mapped to the same RM state in both $\tau_{m,n}$ and $\tau_{m,n+1}$. Eqs (6-8) form the basic scaffolding of our mapping.
Thus far, all constraints have been applied over individual or subsequent tuples of $T_{o}$. However, it is necessary to form a consensus on the mapping for all observed traces onto the ARMDP to ensure Property 3 and Property 4. We begin with Property 3.
Property 3 encodes a transition function, $\tilde{T}$, constructed over components $T$, from the NMRDP, and $\delta_{u}$, from the RM. Even though $\delta_{u}$ is unknown, we know that it must be deterministic. That is, for any transition, $(s,a,s^{\prime})\in S\times A\times S$, and for any $u_{i}\in U$, $\delta_{u}(u_{i},L(s,a,s^{\prime}))$ must map to a single $u_{j}\in U$. Therefore, we must constrain $O$ to agree on the assignments of $u_{i}$ and $u_{j}$ over shared transition tuples. We construct the following indicator variable:
$I_{s,a,s^{\prime},i,j}=\left[\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1,\\ (s,a,s^{\prime})=(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\end{subarray}}^{l_{m}}O_{m,n,i,j}\geq 1\right]$
$I_{s,a,s^{\prime},i,j}$ is an indicator variable over the sum of all $O_{m,n,i,j}$ in $O$ with transition tuple, $(s,a,s^{\prime})$, and assignments $i,j\in\{1,2,...,K\}$. If $I_{s,a,s^{\prime},i,j}=1$, then our resulting ARMDP encodes the transition $((s,u_{i}),a,(s^{\prime},u_{j}))$ in $\tilde{T}$, and thus, $\delta_{u}(u_{j},L(s,a,s^{\prime}))=u_{j}$. If we let $(S,A,S)_{o}=\{(s_{n}^{m},a_{n}^{m},s_{n+1}^{m}),\forall\tau_{m,n}\in\tau_{m},\forall\tau_{m}\in T_{o}\}$ be the set of all unique transitions in $T_{o}$, then to enforce the determinism of $\delta_{u}$ we impose that $\forall(s,a,s^{\prime})\in(S,A,S)_{o},\forall i=1,2,...,K$:
$\sum_{j=1}^{K}I_{s,a,s^{\prime},i,j}\leq 1$ (9)
Consider that for some $(s,a,s^{\prime})\in(S,A,S)_{o}$, and for some fixed $i\in\{1,2,...,K\}$ that when we constrain the sum $\forall j=1,2,...,K$ to be less than or equal to 1, we limit $\delta_{u}(u_{i},L(s,a,s^{\prime}))$ to have at most one outcome. We note that Equation (9) would be an equality when spanning $S\times A\times S$ instead of $(S,A,S)_{o}$. Equation (9), then, satisfies the determinism of $\delta_{u}$ and, therefore, Property 3 of the ARMDP.
The encoding of Property 4 enforces the resolution of reward conflicts in $T_{o}$. While conflicts exist under the perspective of $S\times A\times S$, by mapping conflicting reward tuples onto $S\times U\times A\times S\times U$ we can separate them. Similar to our encoding of Property 3, we do this by constraining $O$ to agree on the assignments of $u_{i}$ and $u_{j}$ over shared reward tuples by way of indicator variables:
$I_{s,a,r,s^{\prime},i,j}=\left[\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1,\\ (s,a,r,s^{\prime})=\tau_{m,n}\end{subarray}}^{l_{m}}O_{m,n,i,j}\geq 1\right]$
$I_{s,a,r,s^{\prime},i,j}$ is an indicator variable over the sum of all $O_{m,n,i,j}$ in $O$ with reward tuple, $(s,a,r,s^{\prime})$, and assignments $i,j\in\{1,2,...,K\}$. If $I_{s,a,r,s^{\prime},i,j}=1$, the ARMDP encodes $\tilde{R}((s,u_{i}),a,(s^{\prime},u_{j}))=r$, which is equivalent to $\delta_{r}(u_{i},L(s,a,s^{\prime}),u_{j})=r$ in the RM. If we let $r(s,a,s^{\prime})=\{r_{n}^{m},\forall\tau_{m,n}\in\tau_{m},\forall\tau_{m}\in T_{o}\mid(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=(s,a,s^{\prime})\}$ be the set of all observed rewards for some $(s,a,s^{\prime})$, then to enforce the determinism of $\delta_{r}$ we impose that $\forall(s,a,s^{\prime})\in(S,A,S)_{o},\forall i=1,2,...,K,\forall j=1,2,...,K$:
$\sum_{r\in r(s,a,s^{\prime})}I_{s,a,r,s^{\prime},i,j}\leq 1$ (10)
Consider that for some $(s,a,s^{\prime})\in(S,A,S)_{o}$ and for some fixed $i,j\in\{1,2,...,K\}$ that when we constrain the sum $\forall r\in r(s,a,s^{\prime})$ to be less than or equal to 1, we limit $\delta_{r}(u_{i},L(s,a,s^{\prime}),u_{j})$ to have at most one outcome, thus resolving our reward conflicts. Importantly, by Equation (10), we encode rewards as conditionals for inferring the structure of the RM. Therefore, if upstream rewards form dependencies with downstream rewards, rewards serve as memory, hence $H=(S\times A\times\mathbb{R})$. Like in Equation (9), Equation (10) would be an equality when spanning $\mathbb{R}$ and $S\times A\times S$ instead of some finite observed space. Regardless, by Equation (10) we satisfy the determinism of $\delta_{r}$ and, thereby, Property 4 of the ARMDP.
Eqs (6-10) form the entirety of the constraints for ARMDP learning. However, there are potentially many ARMDP solutions for some given $T_{o}$. As such, we are motivated to derive simple and interpretable models for representing reward dynamics. Therefore, we orient our objective and learning procedure towards finding the most minimal ARMDP, and thus the most minimal RM, that satisfies $T_{o}$. This is done in two ways. First, for any inference of an ARMDP, we start by assuming $K=|U|=2$, thus dramatically limiting our search space. If $K=2$ is determined to be insufficient to resolve reward conflicts (ILP infeasibility) we increment $K$ by 1 and try again. This is known as a deepening strategy employed in various related works [9, 6]. Second, we limit the number of transition tuples that trigger transitions between RM states (e.g., $\delta_{u}(u_{i},L(s,a,s^{\prime}))=u_{j}$ where $u_{i}\neq u_{j}$). We do so by minimizing the following objective:
$z=\sum_{(s,a,s)\in(S,A,S)_{o}}\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1,\\ i\neq j\end{subarray}}^{K}I_{s,a,s^{\prime},i,j}$ (11)
We solve Eqs (6-11) using out-of-the-box ILP solvers in Gurobi [14] to infer viable ARMDPs. We note that all notation in this section was constructed assuming $L:S\times A\times S\rightarrow 2^{P}$ though it is straightforward to rewrite it as being sensitive to $L:S\rightarrow 2^{P}$. We provide the full ILP in Appendix A.
We also note a structural assumption that we make in later experimentation. That is, we only use $o_{m,n,i,j}$ where $j\geq i$ (i.e., the upper triangle matrix of each $O_{m,n}$). The implication of this is that we assume the RM to never return to a previous state, once left. While this might seem to limit the expressiveness of any resulting ARMDP we note that our ILP simply unrolls these cycles, under this assumption, and expands $U$ as needed. Because $T_{o}$ represents a finite set of finite trajectories, learned ARMDPs will be equivalently expressive, but larger in some cases. However, for all RMs considered in this work (e.g., the RMs in Figure 1) this structural assumption holds and we still learn the minimal ARMDP while dramatically reducing learning time. We provide more details on this assumption in Appendix B.
4.3 Extracting RMs from ARMDPs
We provide a brief outline for extracting an RM from an ARMDP. Given that $K=|U|$ is determined by our deepening strategy, we assume $U$ is given. Additionally, by Equation (7) we designate $u_{1}\in U$, to be the initial state. To extract the rules of $\delta_{u}$ and $\delta_{r}$ we simply iterate $\forall m=1,2,...M,\forall n=1,2,...,l_{m},\forall i=1,2,...,K,\forall j=1,2,...,K$. If $O_{m,n,i,j}=1$ then $\delta_{u}(u_{i},L(s_{n}^{m},a_{n}^{m},s_{n+1}^{m}))=u_{j}$ and $\delta_{r}(u_{i},L(s_{n}^{m},a_{n}^{m},s_{n+1}^{m}),u_{j})=r_{n+1}^{m}$.
4.4 Theoretical Considerations
Our ILP is constructed such that it infers an ARMDP that satisfies reward conflicts in $T_{o}$. However, because the ARMDP is simply just an MDP, we can apply out-of-the-box RL algorithms to solve it. However, it remains to be shown how the ARMDP might be used as a proxy of the NMRDP for the purpose of maximizing reward expectations. Because we assume the form of $R$ in the NMRDP to be an RM, then any observed reward in $T_{o}$ would be consistently reproducible assuming identical histories. Therefore, so long as the likelihood of producing trajectories under $T$ is identical to $\tilde{T}$, reward expectations are preserved. We start with the weighted reward sum for some $\tau_{m}\in T_{o}$:
$R(\tau_{m})=\sum_{n=1}^{l_{m}}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})$ (12)
Similarly, we construct a weighted reward sum for some $\tilde{\tau}_{m}$ derived by mapping $\tau_{m}$ onto an ARMDP to be:
$R(\tilde{\tau}_{m})=\sum_{n=1}^{l_{m}}\sum_{i=1}^{K}\sum_{j=1}^{K}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\cdot O_{m,n,i,j}$ (13)
We remind readers that every $O_{m,n,i,j}$ reflects a binary variable that, if 1, allows us to conclude $\delta_{u}(u_{i},L(s_{n}^{m},a_{n}^{m},s_{n+1}^{m}))=u_{j}$. By referencing Property 3 of the ARMDP, the interpretation of multiplying $O_{m,n,i,j}$ into $T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})$ to determine $\tilde{T}$ should be evident. With this we introduce Theorem 1:
Theorem 1.
$R(\tau_{m})=R(\tilde{\tau}_{m})$ $\forall\tau_{m}\in T_{o}$ where $\tilde{\tau}_{m}$ is derived by mapping $\tau_{m}$ onto an ARMDP.
For space reasons, we provide the proof of Theorem 1 in Appendix C, but state that by Theorem 1, the ARMDP is a suitable proxy for maximizing reward sums $\forall\tau_{m}\in T_{o}$.
4.5 RL for Non-Markov Rewards
Our ILP formulation for learning ARMDPs is a form of passive learning. That is to say, the ARMDP will be sufficient for $T_{o}$, but may not reflect the ground truth. However, by Theorem 1, we know that we can use ARMDPs as a proxy for maximizing reward expectations for the NMRDP. Therefore, we apply Q-Learning to make our approach an active learner. Our intuition is that we might resolve the reward conflicts that are needed to maximize our expected reward.
Algorithm 1 ARMDPQ-Learning
Input: NMRDP
1:  $ARMDP=ARMDP_{0}$
2:  $\pi_{ARMDP}=\pi_{0}$
3:  $T_{o}=[]$
4:  $K=|U|=2$
5:  while $!done$ do
6:     $\tau_{m},conflict=NMRDP.sim(ARMDP,\pi_{ARMDP})$
7:     if $conflict$ then
8:        $T_{o}.append(\tau_{m})$
9:        $solved=False$
10:        while $!solved$ do
11:           $ARMDP,solved=solve\_ILP(T_{o},K)$
12:           if $!solved$ then
13:              $K=K+1$
14:           end if
15:        end while
16:        $\pi_{ARMDP}=\pi_{0}$
17:     else
18:        $\pi_{ARMDP}=update\_qsa(\tau_{m})$
19:     end if
20:  end while
We present ARMDPQ-Learning, with pseudocode available in Algorithm 1. The agent applies our deepening strategy by initially assuming the problem is an MDP (e.g., $|U|=1$). We denote this ARMDP as ARMDP0 (line 1). $K=|U|=2$, is used for subsequent ARMDP inference. The agent policy, $\pi_{ARMDP}$, is set to $\pi_{0}$, a policy with all Q-values set to 0 (line 2). We use the NMRDP to simulate some $\tau_{m}$ under the perspective of ARMDP, selecting actions according to $\pi_{ARMDP}$ (line 6). If a conflict is found we add $\tau_{m}$ into $T_{o}$ (line 8) and solve for a new ARMDP (line 11). If the problem is infeasible we increment $K$ by 1 (line 13) until there is a solution. As each ARMDP presents a new perspective, we reset $\pi_{ARMDP}$ back to $\pi_{0}$ (line 16). Assuming no conflict is found for $\tau_{m}$, we update $\pi_{ARMDP}$ with $\tau_{m}$ according to Q-Learning updating (line 18). Note that line 18 is shorthand. In reality, tuples are updated in an online fashion while simulating the trajectory. Also, by only adding $\tau_{m}$ known to conflict with our current ARMDP, we keep $T_{o}$ small to make our ILP solving more efficient.
5 Experiments
We briefly outline our experiments. Experiment 1 tests the efficacy of ARMDPQ-Learning in the Officeworld domain. Experiment 2 assesses the representative power of ARMDPs by comparing learning profiles of DQNs using $\tilde{S}$ versus various RDQN architectures only using $S$. Finally, Experiment 3 demonstrates the efficacy of learning RMs over DFAs given their ability to learn over $H=(S\times A\times\mathbb{R})^{*}$.
5.1 Experiment 1 - Officeworld
We validate our approach by applying ARMDPQ-Learning on the four black-box RMs (Tasks (b-e)) from the Officeworld domain visualized in Figure 1. We extend each RM with a terminal state on all * tiles, resulting in a Markov reward of -1. The RM, as well as $P$ and $L$ are hidden from the agent. We refer readers to Appendix D for details on hyper-parameter selection, but note that we use an $\epsilon$-greedy policy and set $\epsilon$ to 0% after 90,000 episodes to exploit learned behavior. Results for Experiment 1 can be seen in Figure 2.
Figure 2: ARMDPQ-Learning evaluated on Tasks (b-e) in the Officeworld domain on 10 random trials. Rewards show +/- a standard deviation.
The agent learns to perform optimally in all tasks, suggesting that a meaningful representation was learned. Task (b), (c), and (e) all arrived at $|U|$ reflecting the ground truth model (2, 2, and 4, respectively). Conversely, Task (d) arrived at $|U|\approx 3.2$. Because the coffee and the mail tiles can be achieved in any order, occasionally the agent would lock on to one solution, and not the other, requiring less RM states. $T_{o}$ remained relatively small for all tasks, with $|T_{o}|\approx$ 4, 8, 39, and 44 for Tasks (b-e), respectively. We also tracked the cumulative solve times for each task to be $\approx$ 0.1, 0.6, 4,200, and 12,000 seconds, respectively. We include an extension to Figure 2 in Appendix D, containing figures for $|U|$, $|T_{o}|$ and cumulative solve times.
In all experiments, we noticed high variability in early performance reflected in high standard deviations between episodes 0-20,000. We attribute this to identifying the wrong pattern in $H$ during early conflict resolution. Interestingly, because these early solutions are so poor, they are easily testable by the Q-Learing agent. This lends some credence to our initial intuition of the Q-Learning agent serving as an important active learner capable of interrogating its own perspectives and updating in the face of new evidence.
5.2 Experiment 2 - ARMDP Representations
Previously, Christoffersen et al. compared the learning rates of Q-Learning agents over NMRDPs, using DFAs, to sparse, recurrent deep RL models [9]. However, in smaller discrete state spaces, Q-Learning has the advantage of updating state-action values directly given their simpler, tabular representation. In contrast, deep RL models must learn approximate solutions using complex computations over large batches of data. It remains to be shown how beneficial it could be to represent the state space of a DQN with the state space of an ARMDP. This would allow for a more even comparison in learning performance against RDQN architectures that have to learn a Markov representation over the sparse histories they encode.
For each Task (b-e), we extract the ARMDP representations from Experiment 1 and represent $S$ and $U$ as the one-hot encoded feature space of an Abstracted DQN (ADQN). As the ADQN explores each task we update $S$ and $U$ according to its respective ARMDP. In parallel, we train recurrency-based models without access to these abstractions, namely, a Recurrent DQN (RDQN), a Long-Short Term Memory DQN (LSTMDQN), and a Gated Recurrent DQN (GRUDQN). For a baseline, we also train a DQN without any knowledge of $U$.
Figure 3: DQN models evaluated on Tasks (b-e) from the Officeworld domain on 10 random trials.
Rewards are very sparse for the Officeworld tasks and exploration via an $\epsilon$-greedy policy can often be insufficient. As we don’t seek to demonstrate the efficacy of RL strategies under sparse rewards, we occasionally seed each model with an optimal trajectory (every 10th trajectory). Results for Experiment 2 can be seen in Figure 3.
In all tasks, the ADQN model learns at the rate of, or out-learns, all other models. This becomes more evident as the complexity of the task increases (Task (e) being the most complex). We note that for Task (b), the baseline DQN performs just as well as the other models only because the shortest path to the goal includes the coffee tile. Interestingly, for Task (e), the RDQN performs just as poorly as the baseline DQN model while the LSTMDQN and GRUDQN models (while learning slowly) appear to be uncovering the relevant hidden mechanisms of the underlying RM. We hypothesize that the forgetting component of the GRUDQN and LSTDQN is a key differentiating factor in their learning ability. Importantly, we remind readers to still consider the cumulative solve times of ARMDQ-Learning from Experiment 1 as this is a non-trivial time cost.
5.3 Experiment 3 - Breakfastworld
To evaluate the effectiveness of learning our RMs over $H=(S\times A\times\mathbb{R})^{*}$, rather than $H=(S\times A)^{*}$, we introduce a new domain, Breakfastworld, seen in Figure 4 (a). We include two tasks, Tasks (b) and (c), containing multiple interdependent reward signals. The domain uses $P=\{c,w,e,l\}$ representing propositions for cooking, washing, eating and leaving, respectively. Task (b) contains an RM wherein the agent is meant to cook, eat, and then leave. $(u_{1},c,u_{2})$ returns a reward of -0.1, reflecting a time cost. However, this time cost is experienced in a non-linear fashion as the agent grows bored and spending another turn cooking returns a reward of -0.3, $(u_{2},c,u_{3})$. However, this patience pays off in $(u_{3},e,u_{4})$, where eating a perfectly cooked meal returns a reward of 2. In contrast, $(u_{2},e,u_{5})$ returns 1, reflecting an undercooked meal. Finally, $(u_{4},l,term)$ and $(u_{5},l,term)$ reflect the last reward dependency, returning the cumulative reward for each path in the RM. Task (c) builds on Task (b) by adding a reward dependency for washing. The penalty for washing after cooking once, $(u_{6},w,u_{7})$, is less than after cooking twice, $(u_{4},w,u_{5})$.
Figure 4: The Breakfastworld domain (a) with two RMs, (b) to cook, eat and then leave, (c) to cook, eat, wash, and then leave.
Both tasks reflect RMs more complex in terms of $|U|$ than Task (e) in Experiment 1. However, we will demonstrate that RM learning benefits in cases where rewards are interdependent, as observed reward conflicts result in natural splits in the ARMDP learning by Equation (10). Conversely, DFA learning suffers under such conditions as each reward conflict must be represented by its own DFA [13]. Moreover, downstream DFAs are learned independently from upstream DFAs as they only learn according to $H=(S\times A)^{*}$.
We demonstrate how advantageous it can be that our ILP naturally learns according to $H=(S\times A\times\mathbb{R})^{*}$ by applying ARMDPQ-Learning to Tasks (b) and (c). We then compare this to applying ARMDPQ-Learning on RMs equivalent in structure to Task (b) and (c), but that only emit the cumulative rewards on the terminal state. All other rewards are made 0. We refer to Tasks (b) and (c), in Figure 4, as Task (b) Full and Task (c) Full, respectively. Conversely, the RMs that only emit the cumulative reward are referred to as Task (b) Cumulative and Task (c) Cumulative, respectively. The purpose of these cumulative tasks is to form a problem that produces reward conflicts on tile [0,0] that can not be contextualized by the prior reward conflicts. This would be similar to learning two DFAs to parse the two different possible rewards on tile [0,0] only using $H=(S\times A)^{*}$. We show results for Experiment 3 in Table 1.
In all solvable cases of ARMDP-Learning, the agent achieved optimal performance and identified the correct $|U|$ for each task. We include an extension figure to Table 1 in Appendix F containing such results. Using Task (b) Full our cumulative solve time ($\approx$ 4.4 seconds) was dramatically faster than when using Task (b) Cumulative ($\approx$ 1,102.9 seconds). This can be attributed to Task (b) Full requiring roughly a fourth of the number of trajectories to solve than Task (b) Cumulative, likely because of the leveraged context of upstream rewards. More impressive were the results for Task (c) Full which resulted in the cumulative solve time of $\approx$ 153.8 seconds. In no instance of attempting to learn in Task (c) Cumulative did we arrive at a stable solution in 100,000 seconds ($>$27 hours).
Task $|T_{o}|$ Solve Time (s)
Task (b) Full 6.2 $\pm$ 1.1 4.4 $\pm$ 3.3
Task (b) Cumulative 22.4 $\pm$ 8.8 1102.9 $\pm$ 1068.8
Task (c) Full 10.8 $\pm$ 3.0 153.8 $\pm$ 197.8
Task (c) Cumulative n/a n/a
Table 1: Results of ARMDPQ-Learning evaluated on Tasks (b) and (c) from the Breakfastworld domain on 10 random trials.
6 Conclusions
In this work, we introduced a formulation for learning RMs from non-Markov rewards by solving an ILP to learn an ARMDP. The ARMDP represents the cross-product between an RM and the underlying state space of an NMRDP. Importantly, we learn the ARMDP, and thus the RM, without knowledge of $P$, by identifying hidden triggers that govern the rules of $L(S\times A\times S)\rightarrow 2^{P}$. We also demonstrated how effective it can be to learn our automaton according to $H=(S\times A\times\mathbb{R})^{*}$, rather than $H=(S\times A)^{*}$, a limitation of DFAs. We proved that the ARMDP shares an equivalent expectation to the NMRDP for $T_{o}$ and can therefore be used as a proxy for maximizing reward expectations. We validated our approach by applying ARMDPQ-Learning to train Q-Learning agents over black-box RMs. We also evaluated the representative power of the ARMDP by representing the state space of a DQN as the state space of an ARMDP. We compared the training profiles of these DQNs to the training profiles of various RDQN architectures that had to learn Markov representations over the sparse histories they encode.
One important aspect of this work is that the RMs we learn are inherently more interpretable than the sparse representation of histories in $H=(S\times A\times\mathbb{R})^{*}$. This is in part because we learn minimal RMs that encode $H$ under the finite, structured, and symbolic representation of an automaton. However, more importantly, interpreting an RM doesn’t require a comprehensive understanding of $S$ and $A$ outside of what the labeler encodes. This is a consequence of the RM being entirely decomposable from the ARMDP. For example, the RM in Figure 1 Task (b) conveys that we need to bring coffee to the office without ever needing to describe where the coffee might be sourced from or where the office is. Since RL is predicated on the assumption that the reward function is the most concise, robust, and transferable definition of a task [20], we posit that RMs emerge as strong candidates for reward function representation.
RM representations also carry important weight for the field of Inverse RL (IRL). In IRL, the goal is to learn a reward function from behavioral data that motivated the behavior [20, 26, 8]. While most approaches assume a Markov reward function, some of this work has been focused on learning rewards in Partially Observable MDPs (POMDPs) [8] and in Semi MDPs (SMDPs) [26]. We think that relaxing the Markov assumption in a principled way, analogous to this work, could yield insight into hidden abstractions made by a decision-maker. Moreover, given their interpretable form, it might shed light on what these hidden abstractions represent.
Lastly, we generally make note that this work is limited in that it assumes $S$ and $A$ are discrete. As such, we would like to consider automata extraction methods, in the future, using RNNs [28, 21, 31] as this would extend our approach into the continuous case. Weiss et al. have demonstrated promising results in this regard by applying $L*$ as an exact learning algorithm for deep recurrent models [31].
Appendix
Appendix A Full ILP
$min\hskip 2.0ptz=$ $\sum_{(s,a,s)\in(S,A,S)_{o}}\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1,\\ i\neq j\end{subarray}}^{K}I_{s,a,s^{\prime},i,j}$ $\hskip 2.0pt(1)$
$s.t.$
$\sum_{i=1}^{K}\sum_{j=1}^{K}O_{m,n,i,j}=1$ $\forall m=1,2,...,M,\forall n=1,2,...l_{m}$ $\hskip 2.0pt(2)$
$\sum_{j=1}^{K}O_{m,1,1,j}=1$ $\forall m=1,2,...,M$ $\hskip 2.0pt(3)$
$\sum_{i=1}^{K}O_{m,n,i,j}=\sum_{i=1}^{K}O_{m,n+1,j,i}$ $\forall m=1,2,...,M,\forall n=1,2,...,l_{m}-1,\forall j=1,2,...,K$ $\hskip 2.0pt(4)$
$\sum_{j=1}^{K}I_{s,a,s^{\prime},i,j}\leq 1$ $\forall(s,a,s^{\prime})\in(S,A,S)_{o},\forall i=1,2,...,K$ $\hskip 2.0pt(5)$
$\sum_{r\in r(s,a,s^{\prime})}I_{s,a,r,s^{\prime},i,j}\leq 1$ $\forall(s,a,s^{\prime})\in(S,A,S)_{o},\forall i=1,2,...,K,\forall j=1,2,...,K$ $\hskip 2.0pt(6)$
$I_{s,a,s^{\prime},i,j}=\left[\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1,\\ (s,a,s^{\prime})=(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\end{subarray}}^{l_{m}}O_{m,n,i,j}\geq 1\right]$ $\forall(s,a,s^{\prime})\in(S,A,S)_{o},\forall i=1,2,...,K,\forall j=1,2,...,K$ $\hskip 2.0pt(7)$
$I_{s,a,r,s^{\prime},i,j}=\left[\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1,\\ (s,a,r,s^{\prime})=\tau_{m,n}\end{subarray}}^{l_{m}}O_{m,n,i,j}\geq 1\right]$ ${\LARGE\begin{subarray}{c}\forall(s,a,s^{\prime})\in(S,A,S)_{o},\forall r\in r(s,a,s^{\prime}),\\ \hskip 30.0pt\forall i=1,2,...,K,\forall j=1,2,...,K\end{subarray}}$ $\hskip 2.0pt(8)$
$O_{m,n,ij}\in\{0,1\}$ ${\LARGE\begin{subarray}{c}\forall m=1,2,...,M,\forall n=1,2,...,l_{m},\\ \hskip 13.0pt\forall i=1,2,...,K,\forall j=1,2,...,K\end{subarray}}$ $\hskip 2.0pt(9)$
The goal of our Integer Linear Program (ILP) formulation is to infer a valid Abstract Reward Markov Decision Process (ARMDP), as defined in Definition 1, from a set of observed trajectories with observed reward conflicts. Reward conflicts are a result of an agent observing rewards under the lens of $S\times A\times S$ when in reality rewards are sensitive to histories of the form $H=(S\times A\times\mathbb{R})^{*}$. We introduce some notation that our ILP leverages. We derive candidate ARMDPs by solving an ILP over a set of observed trajectory traces, $T_{o}$. Let, $\tau_{m}\in T_{o}$, be the m-th trace in $T_{o}$ of length $l_{m}$. Further, let $\tau_{m,n}=(s_{n}^{m},a_{n}^{m},r_{n+1}^{m},s_{n+1}^{m})$ be the n-th tuple of $\tau_{m}$. We define a reward conflict to be any two $\tau_{m,n}$ and $\tau_{\tilde{m},\tilde{n}}$ where $(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=(s_{\tilde{n}}^{\tilde{m}},a_{\tilde{n}}^{\tilde{m}},s_{\tilde{n}+1}^{\tilde{m}})$, but $r_{n+1}^{m}\neq r_{\tilde{n}+1}^{\tilde{m}}$.
We define the matrix, $O=M\times N\times K\times K$, where $M=|T_{o}|$, $N$ reflects the largest $l_{m}$, $\forall\tau_{m}\in T_{o}$, and $K=|U|$ reflects the size of the finite state space for some candidate RM. $O_{m,n}$, then, is a $K\times K$ matrix that reflects the mapping of $\tau_{m,n}=(s_{n}^{m},a_{n}^{m},r_{n+1}^{m}s_{n+1}^{M})$ onto the ARMDP, where both $s_{n}^{m}$ and $s_{n+1}^{m}$ are assigned an RM state. Additionally, we make use of two sets to gather necessary variables from $T_{o}$ for the construction of our ILP. These are $(S,A,S)_{o}=\{(s_{n}^{m},a_{n}^{m},s_{n+1}^{m}),\forall\tau_{m,n}\in\tau_{m},\forall\tau_{m}\in T_{o}\}$ and $r(s,a,s^{\prime})=\{r_{n}^{m},\forall\tau_{m,n}\in\tau_{m},\forall\tau_{m}\in T_{o}\mid(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=(s,a,s^{\prime})\}$ where $(S,A,S)_{o}$ reflects the set of all observed transition tuples in $T_{o}$ and $r(s,a,s^{\prime})$ is the set of all observed rewards for some observed transition tuple in $T_{o}$.
Appendix B j $\geq$ i Constraint
This section describes the implications of imposing a $j\geq i$ constraint in the construction of $O$. Typically, $O_{m,n}$ is a $K\times K$ matrix that reflects the mapping of $\tau_{m,n}=(s_{n}^{m},a_{n}^{m},r_{n+1}^{m}s_{n+1}^{M})$ onto the ARMDP, where both $s_{n}^{m}$ and $s_{n+1}^{m}$ are assigned an RM state:
$O_{m,n}=\begin{pmatrix}O_{m,n,1,1}&O_{m,n,1,2}&\cdots&O_{m,n,1,K}\\ O_{m,n,2,1}&O_{m,n,2,2}&\cdots&O_{m,n,2,K}\\ \vdots&\vdots&\ddots&\vdots\\ O_{m,n,K,1}&O_{m,n,K,2}&\cdots&O_{m,n,K,K}\end{pmatrix}$
When we impose a $j\geq i$, we only consider the upper right triangle matrix of $O_{m,n}$:
$O_{m,n}=\begin{pmatrix}O_{m,n,1,1}&O_{m,n,1,2}&\cdots&O_{m,n,1,K}\\ &O_{m,n,2,2}&\cdots&O_{m,n,2,K}\\ &&\ddots&\vdots\\ &&&O_{m,n,K,K}\end{pmatrix}$
The implication of this construction on the form of inferred ARMDP is that the underlying RM can never return to a prior state, once left. To reiterate, this does not limit the expressivity of the inferred ARMDP, it simply requires larger $U$ in some instances. Consider the RMs in Figure 5. Task (a) represents a patrol task identical to Task (e) in Figure 1 in the main paper. Even though Task (a) has atomic cycles (e.g., $u_{1}\rightarrow u_{1}$), once a state is left, the RM never returns to it. For RMs like this, even with $j\geq i$, we will still be able to represent the RM minimally. Conversely, Task (b) is a continuous patrol task, where on completion of a lap, the RM begins again in $u_{1}$. Here, because the RM reuses previous states, the ARMDPs that we produce will continuously unroll these states and expand $U$ as needed. Instead of $u_{1}\rightarrow u_{2}\rightarrow u_{3}\rightarrow u_{4}\rightarrow u_{1}...$, the RM we infer will be represented as $u_{1}\rightarrow u_{2}\rightarrow u_{3}\rightarrow u_{4}\rightarrow u_{5}...\hskip 2.0pt$. However, because $T_{o}$ represents a finite set of finite trajectories, we will always be able to learn an ARMDP from it, though perhaps at the expense of excess computation in cycle decomposition. Regardless, all inferred ARMDPs will be equally expressive.
Figure 5: Patrol task where (a) patrols only once and (b) patrols continuously
Appendix C ARMDPs as NMRDP Proxies
The goal of this section is to prove that the ARMDP is a sufficient proxy of the NMRDP for the purposes of maximizing reward expectations. Because we assume the form of $R$ in the NMRDP to be an RM, then any observed reward in $T_{o}$ would be consistently reproducible assuming identical histories. Therefore, so long as the likelihood of producing trajectories under $T$ is identical to $\tilde{T}$, reward expectations are preserved. We restate the the weighted reward sum for some $\tau_{m}\in T_{o}$:
$R(\tau_{m})=\sum_{n=1}^{l_{m}}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})$ (14)
Similarly, we restate the weighted reward sum for some $\tilde{\tau}_{m}$ derived by mapping $\tau_{m}$ onto an ARMDP to be:
$R(\tilde{\tau}_{m})=\sum_{n=1}^{l_{m}}\sum_{i=1}^{K}\sum_{j=1}^{K}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\cdot O_{m,n,i,j}$ (15)
We remind readers that every $O_{m,n,i,j}$ reflects a binary variable that, if 1, allows us to conclude $\delta_{u}(u_{i},L(s_{n}^{m},a_{n}^{m},s_{n+1}^{m}))=u_{j}$. We reference Property 3 of the ARMDP:
$\tilde{T}((s,u),a,(s^{\prime},u^{\prime}))=\begin{cases}T(s,a,s^{\prime})&\text{if\ }\delta_{u}(u,L(s,a,s^{\prime}))=u^{\prime}\\ 0&\text{otherwise}\end{cases}$
By multiplying $O_{m,n,i,j}$ into $T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})$, $O_{m,n,i,j}$ serves as a logical gate to produce the dynamics of $\tilde{T}$. With this we introduce Theorem 1:
Theorem 2.
$R(\tau_{m})=R(\tilde{\tau}_{m})$ $\forall\tau_{m}\in T_{o}$ where $\tilde{\tau}_{m}$ is derived by mapping $\tau_{m}$ onto an ARMDP.
Proof.
$R(\tilde{\tau}_{m})$ $=\sum_{n=1}^{l_{m}}\sum_{i=1}^{K}\sum_{j=1}^{K}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\cdot O_{m,n,i,j}$
$=\sum_{n=1}^{l_{m}}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\sum_{i=1}^{K}\sum_{j=1}^{K}\cdot O_{m,n,i,j}$
$=\sum_{n=1}^{l_{m}}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\cdot 1$
$=\sum_{n=1}^{l_{m}}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=R(\tau_{m})$
Appendix D Experiment 1 Specifications
For all tasks in Experiment 1, we used $L:S\rightarrow 2^{P}$ and $j\geq i$ toggle construction. Hyper parameters include $\gamma=0.95$, $\alpha=0.1$ and $\epsilon=0.1$. The only exception is for Task (b) where we use $\epsilon=0.2$ until the first conflict, to promote exploring. This is because the shortest path to the goal state includes the necessary tile to trigger the non-Markov reward. As such, the Markov representation under $S$ and $A$ appears sufficient to demonstrate optimal behavior. Even with $\epsilon=0.2$, for Task (b), 100,000 episodes was oftentimes not enough to trigger a conflict. To further accommodate this we put a sliding window on the total episode counts used to produce Figure 2 in the main paper such that the 100,000 episodes contained the first conflict. This was strictly to produce a consistent figure over 10 randomly seeded trials. We include an extended version of Figure 2 from the main paper in Figure 6 containing information on $|U|$, $|T_{o}|$, as well as Cumulative Solve Times (in seconds). Machine specifications used in the experiment: 32 core Intel(R) Xeon(R) CPU E5-2640 v3 @ 2.60 GHz with 512 GB of Memory.
Figure 6: Results of ARMDPQ-Learning evaluated on Tasks (b-e) from the Officeworld domain on 10 random trials. Episode Reward Avg column shows average reward +/- a stddev. $|U|$ column shows current assumed size of $U$. $|T_{o}|$ column shows current size of $T_{o}$. Cumulative Solve Time(s) represents the total solve time for all ARMDP inferences, in seconds.
Appendix E Experiment 2 Specifications
Each model in Experiment 2 was trained using a Double Deep Q-Network (DDQN) architecture, meaning each model had a target and non-target model supporting the Deep Q-Network (DQN) architecture. All target models were trained slowly off of their respective non-target model using a smoothing factor of 0.01. Each model across the types (ADQN, DQN, RDQN, GRUDQN, and LSTMDQN) were all trained using a learning rate of 1e-3 with Adam optimizer. ADQN and DQN used a batch size of 32, whereas the RDQN, GRUDQN, and LSTMDQN models used a batch update by stacking a single trajectory. Other hyper parameters include $\gamma=0.95$ and $\epsilon=0.1$. Machine specifications used in the experiment: 32 core Intel(R) Xeon(R) CPU E5-2640 v3 @ 2.60 GHz with 512 GB of Memory.
Appendix F Experiment 3 Specifications
For all tasks in Experiment 3, we used $L:S\rightarrow 2^{P}$ and $j\geq i$ toggle construction. Hyper parameters include $\gamma=0.95$, $\alpha=0.1$ and $\epsilon=0.1$. We applied a similar sliding window as we did in Experiment 1 but under a lens of 50,000 episodes. We included an extended version of Table 1 in Figure 7. Machine specifications used in the experiment: 32 core Intel(R) Xeon(R) CPU E5-2640 v3 @ 2.60 GHz with 512 GB of Memory.
Figure 7: Results of ARMDPQ-Learning evaluated on Tasks (b) and (c) from the Breakfastworld domain on 10 random trials. Task (b) and (c) Full show results when using the full RM for Task (b) and (c). Task (b) Cumulative shows results for using the RM for Task(b) where all rewards are 0 except the terminal cumulative rewards. Episode Reward Avg shows average reward +/- a stddev. $|U|$ shows current assumed size of $U$. $|T_{o}|$ shows current size of $T_{o}$. Cumulative Solve Time(s) shows the total solve time, for all ARMDPs, in seconds.

### v3

@@
 Many Reinforcement Learning algorithms assume a Markov reward function to guarantee optimality.
-However, not all reward functions are known to be Markov.
-This paper proposes a framework for mapping non-Markov reward functions into equivalent Markov ones by learning a Reward Machine - a specialized reward automaton.
+However, not all reward functions are Markov.
+This paper proposes a framework for mapping non-Markov reward functions into equivalent Markov ones by learning specialized reward automata, Reward Machines.
 Unlike the general practice of learning Reward Machines, we do not require a set of high-level propositional symbols from which to learn.
-Rather, we learn hidden triggers directly from data that encode them.
-Given their ability to model reward dependencies, we demonstrate the importance of learning Reward Machines versus their Deterministic Finite-State Automata counterparts.
+Rather, we learn hidden triggers, directly from data, that construct them.
+We demonstrate the importance of learning Reward Machines over their Deterministic Finite-State Automata counterparts given their ability to model reward dependencies.
 We formalize this distinction in our learning objective.
 Our mapping process is constructed as an Integer Linear Programming problem.
-We prove that our mappings provide consistent expectations for the underlying process.
-We empirically validate our approach by learning black-box non-Markov Reward functions in the Officeworld Domain.
-Additionally, we demonstrate the effectiveness of learning dependencies between rewards in a new domain, Breakfastworld.
+We prove that our mappings form a suitable proxy for maximizing reward expectations.
+We empirically validate our approach by learning black-box, non-Markov reward functions in the Officeworld domain.
+Additionally, we demonstrate the effectiveness of learning reward dependencies in a new domain, Breakfastworld.
+[…]
+2377
 1 Introduction
@@
 For example, Mnih et al.
-[19] stacked subsequent frames from the Atari games [3] to encode a suitable representation for training Deep Q-Networks.
-Similarly, deep, recurrent models have been used encode state histories for similar purposes [15, 11].
+[18] stacked subsequent frames from the Atari games [3] to encode a suitable representation for training Deep Q-Networks (DQNs).
+Similarly, Recurrent DQNs (RDQNs) have been used to encode state histories for similar purposes [14].
 However, these sparse, deep models obscure the interpretation of rewards.
-To better understand them it is crucial to examine their dynamics under less sparse and more interpretable forms.
-A common strategy to represent non-Markov rewards in an interpretable format is to use the finite states of an automata to represent hidden or historical features that render the representation Markov.
-Historically, this has been accomplished with Deterministic Finite-State Automata (DFA) [5, 13, 12], but more recently, specialized reward automata, Reward Machines (RMs), have been used [32, 22, 23, 10].
-We highlight two important distinctions between DFAs and RMs: 1.) DFAs are inherently less expressive than RMs as they only accept or reject traces.
+To better understand them, it is crucial to examine their dynamics under less sparse and more interpretable forms.
+To this end, automata have been used to compactly represent the hidden or historical features of non-Markov rewards.
+While Deterministic Finite-State Automata (DFAs) have traditionally served this role [5, 12, 11], more recently, specialized reward automata, Reward Machines (RMs), have been used [31, 21, 22, 10].
+We highlight two distinctions between DFAs and RMs: 1.) DFAs are inherently less expressive than RMs as they only accept or reject traces.
 Consequently, when rewards are complex, multiple DFAs must be learned [12].
-In contrast, RMs represent complex reward behavior in a single automaton.
+In contrast, RMs can encode complex reward behavior in a single automaton.
 2.) DFAs are only capable of representing patterns over $H=(S\times A)^{*}$, the history of states and actions.
-Conversely, RMs are capable of representing patterns over $H=(S\times A\times\mathbb{R})^{*}$, the history of states, actions, and rewards.
-This affords us a unique learning objective where rewards serve as memory allowing us to model reward dependencies.
-However, prior work for learning RMs assumes access to $P$, the set of high-level propositional symbols the RM operates over, and $L:S\times A\times S\rightarrow 2^{P}$, a labeling function mapping transitions to instantiations in $P$.
+Conversely, RMs can represent patterns over $H=(S\times A\times\mathbb{R})^{*}$, which include rewards.
+This affords us a unique learning objective where rewards serve as memory allowing us to model their dependencies.
+However, prior work on learning RMs assumes access to $P$, a set of high-level propositions, and $L:S\times A\times S\rightarrow 2^{P}$, a labeling function mapping transitions to instantiations in $P$.
 This makes much of this prior work ungeneralizable.
-We instead infer $L$ by identifying minimal patterns in $H=(S\times A\times\mathbb{R})^{*}$, called hidden triggers, that differentiate reward outcomes.
-Rather than learn the RM directly, we map observed non-Markov rewards onto the cross-product of $U$, the state space of a candidate RM, and $S$, the observed state space, to form a Markov representation called the Abstract Reward Markov Decision Process (ARMDP).
-We infer our ARMDPs, and thus its decomposable RM, by solving an Integer Linear Program (ILP), lending itself to powerful off-the-shelf discrete optimization solvers.
-Contributions
-We introduce a novel algorithm for mapping non-Markov reward functions to equivalent Markov ones.
-We show how leveraging $H=(S\times A\times\mathbb{R})^{*}$ in our learning objective can dramatically expedite learning in cases with multiple interdependent reward signals.
-We validate our approach in several ways.
-First, we demonstrate our ability to map non-Markov rewards sampled from black-box RMs in the Officeworld domain, originally introduced in the first RM paper [16].
-Secondly, we evaluate the representative power of the ARMDP by extending DQN agents with ARMDP features to compare their learning profiles against various sparse Recurrent DQN (RDQN) variants.
+We, instead, infer $L$ by identifying minimal patterns in $H=(S\times A\times\mathbb{R})^{*}$, called hidden triggers, that differentiate reward outcomes.
+Rather than learn RMs directly, we map observed non-Markov rewards onto the cross-product of $U$, the state space of a candidate RM, and $S$, the observed state space, to form a Markov representation called the Abstract Reward Markov Decision Process (ARMDP).
+We infer an ARMDP, and thus its decomposable RM, by solving an Integer Linear Program (ILP), lending itself to powerful off-the-shelf discrete optimization solvers.
+Contributions: We introduce a novel algorithm for mapping non-Markov reward functions to equivalent Markov ones.
+We show how by leveraging $H=(S\times A\times\mathbb{R})^{*}$ in our learning objective we can dramatically expedite learning in cases with interdependent reward signals.
+We validate our approach by demonstrating our ability to learn black-box RMs from the Officeworld domain, originally introduced in the first RM paper [15].
+We then evaluate the representative power of the ARMDP by extending DQN agents with ARMDP features to compare their learning profiles against various sparse RDQN variants.
 Finally, we validate our approach to represent complex reward function behavior in a new domain, Breakfastworld, featuring black-box RMs with interdependent reward signals.
-Importantly, we show how by leveraging $H=(S\times A\times\mathbb{R})^{*}$ we can expedite learning in these cases.
-Our contributions are three-fold:
+We show how by leveraging $H=(S\times A\times\mathbb{R})^{*}$ in these experiments we can expedite learning.
+Our main contributions are three-fold:
   1. 1.
-    We introduce a novel algorithm for RM learning without assumed access to $P$ and $L$,
+    We introduce a novel algorithm for learning RMs without assumed access to $P$ and $L$,
   2. 2.
-    We prove that the ARMDP is a suitable representation for maximizing reward expectations under non-Markov rewards,
+    We prove that the ARMDP is a suitable proxy for maximizing reward expectations under non-Markov rewards,
   3. 3.
@@
 2 Related Work
-Decision-making with non-Markov reward functions is often represented under the Non-Markov Reward Decision Process (NMRDP) [27].
+Decision-making with non-Markov reward functions is often represented as a Non-Markov Reward Decision Process (NMRDP) [26].
 An NMRDP is similar to an MDP except that the reward function, $R:H\rightarrow\mathbb{R}$, maps state and action histories, $H=(S\times A)^{*}$, to reward values.
-However, as $H$ can become unwieldy in large state and action spaces, more succinct representations have been encoded using automata.
+However, as $H$ can become unwieldy due to the curse of history, more succinct representations have been encoded using automata.
 The intuition of this approach lies in the fact that automata are finitely represented and offer a level of interpretation due to their structured form and symbolic representation.
 Traditionally, the automata of choice has been the DFA, yielding promising results in identifying Markov representations for NMRDPs [5, 12, 11].
-However, more recently, specialized reward automata, RMs, have been used to represent NMRDP behavior [16].
+More recently, specialized reward automata, RMs, have also been used [15].
 While DFAs operate over an input alphabet and emit true or false over traces, RMs extend this expressiveness by emitting reward values along each transition.
@@
 A notable limitation of these prior works is that they assume access to a set of high-level symbols (the input language for the automata) and the relationship of the underlying state space to those symbols.
-In RM literature these are denoted $P$ and $L:S\times A\times S\rightarrow 2^{P}$ where $P$ is a set of symbols, separate from the NMRDP, and $L$ is a mapping of transitions from the NMRDP to instantiations in $P$, respectively.
+In RM literature, these are denoted $P$ and $L:S\times A\times S\rightarrow 2^{P}$ where $P$ is a set of symbols, separate from the NMRDP, and $L$ is a mapping of transitions from the NMRDP to instantiations in $P$, respectively.
 A similar assumption is often made for DFA learning, with exceptions being the work of Gaon and Brafman [12] as well as Christoffersen et al.
-[9] who learn symbolic representations from $H:(S\times A)^{*}$ to build their automata.
-Gaon and Brafman learn their DFA using L*, however, their approach was shown to be sample inefficient.
+[9] who learn symbolic representations from $H=(S\times A)^{*}$ to build their automata.
+Gaon and Brafman learn their DFA by applying L*, however, their approach was shown to be sample inefficient.
 Christoffersen et al.
@@
 We improve upon these approaches by, instead, learning an RM without assumed access to $P$ and $L$.
-We show how by learning RMs we can model reward dependencies in a single automata by representing patterns in $H:(S\times A\times\mathbb{R})^{*}$.
+We show how, by learning RMs, we can model reward dependencies in a single automata by representing patterns in $H=(S\times A\times\mathbb{R})^{*}$.
 3 Preliminaries
-3.1 RL and Non-Markov Rewards
+3.1 Reinforcement Learning and Non-Markov Reward
 RL is a type of self-supervised learning in which an agent learns to interact with its environment to maximize its long-term reward expectation [24].
 It is an iterative process in which the agent receives feedback in the form of rewards or punishments to adjust its behavior.
-RL is often represented as a sequential decision-making problem formalized using the MDP.
+RL is formalized using the MDP.
 An MDP is a 6-tuple $(S,A,T,R,\gamma,\rho)$, where $S$ is the state space, $A$ is the action space, $T:S\times A\times S\rightarrow[0,1]$ is a Markov probabilistic transition function, $R:S\times A\times S\rightarrow\mathbb{R}$ is a Markov reward function mapping states, actions and states to rewards, $\gamma\in[0,1)$ is the discount factor and $\rho:S\rightarrow[0,1]$ is the initial state distribution.
-We note that NMRDPs are similar to MDPs in that they share $(S,A,T,R,\rho)$, but $R:H\rightarrow\mathbb{R}$ in NMRDPs.
+Briefly, NMRDPs are similar to MDPs in that they share $(S,A,T,R,\gamma,\rho)$; however, $R:H\rightarrow\mathbb{R}$ maps histories of states and actions to rewards, instead.
 The solution for an MDP is a policy, $\pi:S\rightarrow A$, that maps states to actions.
-The value of a policy for any state, $s\in S$, at time point $t$ is the expected discounted return defined as follows:
-$V_{\pi}(s)=\mathbb{E}\left[\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}\mid s_{t}=s\right]$ (1)
-The optimal policy denoted $\pi_{*}$, is determined by solving for the optimal state value function:
-$V_{*}(s)=\max_{\pi}V^{\pi}(s)$ (2)
+The value of a policy for any state, $s\in S$, at time point $t$ is the expected discounted return, defined as follows:
+$V_{\pi}(s)=\mathbb{E}_{\pi}\left[\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}\mid s_{t}=s\right]$ (1)
+where $\mathbb{E}_{\pi}$ is the expected reward from starting in $s$ and following $\pi$. The optimal policy, denoted $\pi_{*}$, is determined by solving for the optimal state value function:
+$V_{*}(s)=\max_{\pi}V_{\pi}(s)$ (2)
 for all $s\in S$. We can write the optimal state-action value function in terms of the state value function as follows:
-$Q_{*}(s,a)=\mathbb{E}\left[r_{t+1}+\gamma V_{*}(s_{t+1})\mid s_{t}=s,a_{t}=a\right]$ (3)
+$Q_{*}(s,a)=\mathbb{E}_{\pi}\left[r_{t+1}+\gamma V_{*}(s_{t+1})\mid s_{t}=s,a_{t}=a\right]$ (3)
 for all $s\in S$ and $a\in A$. Hence:
@@
 RMs operate over a set of propositional symbols, $P$, representing high-level events, separate from $S$.
-RMs are defined as a 5-tuple, $(U,u_{1},F,\delta_{u},\delta_{r})$, where $U$ is a finite state space, $u_{1}\in U$ is the initial state, $F$ is a set of terminal states where $F\cap U=\emptyset$, $\delta_{u}:U\times 2^{P}\rightarrow U\cup F$ is a transition function and $\delta_{r}:U\times 2^{P}\times U\cup F\rightarrow\mathbb{R}$ is a reward function.
+RMs are defined as a 5-tuple, $(U,u_{1},F,\delta_{u},\delta_{r})$, where $U$ is a finite state space, $u_{1}\in U$ is the initial state, $F$ is a set of terminal states where $F\cap U=\emptyset$, $\delta_{u}:U\times 2^{P}\rightarrow U\cup F$ is a transition function and $\delta_{r}:U\times 2^{P}\times(U\cup F)\rightarrow\mathbb{R}$ is a reward function.
 A labeling function, $L:S\times A\times S\rightarrow 2^{P}$, connects $S$ and $A$ to the RM by mapping transitions to instantiations in $P$.
@@
 $\neg$, $\lor$ and $\land$ are logical operators for negation, or, and and, respectively.
-The cross-product of $U$ and $S$ is Markov, which, in the discrete case, means we can solve these problems optimally.
+The cross-product of an NMRDP and an RM can be used to form a Markov representation for predicting both $R$ and $T$.
 4 Approach
@@
 Problem Statement
-From a set of observed, finite-length trajectories, $T_{o}$, sampled from an NMRDP, $(S,A,T,R)$, we learn a minimal RM $(U,u_{1},\delta_{u},\delta_{r})$ that renders rewards Markov to $S$, $A$ and $U$.
+Given a finite set of finite-length trajectories, $T_{o}$, sampled from an NMRDP, $(S,A,T,R,\gamma,\rho)$, we learn a minimal RM $(U,u_{1},\delta_{u},\delta_{r})$ that renders rewards Markov to $S$, $A$ and $U$.
 While $R$ in the NMRDP is unknown, we assume it can be sampled from and is representable as an RM.
-Thus, hereafter, we will refer to $R$ as a black-box RM.
-Our approach to solving this problem is first to learn an ARMDP, which is simply the MDP cross-product between an RM and an NMRDP, that enables the extraction of the associated RM.
+Hereafter, we will refer to $R$ as a black-box RM.
+Our approach for solving this problem is to learn an ARMDP, which is simply the MDP cross-product between an RM and an NMRDP that enables the extraction of the associated RM.
 Learning of the ARMDP is formulated as an ILP.
-A minimal RM is an RM with the smallest hidden state space $|U|$, and the least number of transition edges between states.
-The following sections are organized as follows.
-We begin by defining the ARMDP as the cross-product space between an RM and an NMRDP in Section 4.1.
+A minimal RM is an RM with the smallest hidden state space $|U|$, and the least number of transition edges between different RM states.
+The following sections are organized as follows: We begin by defining the ARMDP as the cross-product space between an RM and an NMRDP in Section 4.1.
 Section 4.2 describes a passive learning approach for learning ARMDPs from $T_{o}$, by solving an ILP.
@@
 Definition 1.
-Given an NMRDP $(S,A,T,R)$, RM $(U,u_{1},\delta_{u},\delta_{r})$, and labeler, $L:S\times A\times S\rightarrow 2^{P}$, an ARMDP is an MDP $(\tilde{S},A,\tilde{T},\tilde{R},\gamma,\tilde{\rho})$ satisfying:
+Given an NMRDP $(S,A,T,R,\gamma,\rho)$, RM $(U,u_{1},\delta_{u},\delta_{r})$, and labeler, $L:S\times A\times S\rightarrow 2^{P}$, an ARMDP is an MDP, $(\tilde{S},A,\tilde{T},\tilde{R},\gamma,\tilde{\rho})$, satisfying:
   1. 1.
@@
 4.2 Formulation
-We infer candidate ARMDPs by solving an ILP, that resolves reward conflicts over $T_{o}$.
-Let, $\tau_{m}\in T_{o}$, be the m-th trace in $T_{o}$ of length $l_{m}$.
+We infer candidate ARMDPs by solving an ILP that resolves reward conflicts over $T_{o}$.
+Let $\tau_{m}\in T_{o}$ be the m-th trace in $T_{o}$ of length $l_{m}$.
 Further, let $\tau_{m,n}=(s_{n}^{m},a_{n}^{m},r_{n+1}^{m},s_{n+1}^{m})$ be the n-th tuple of $\tau_{m}$.
-We define a reward conflict to be any two $\tau_{m,n}$ and $\tau_{\tilde{m},\tilde{n}}$ where $(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=(s_{\tilde{n}}^{\tilde{m}},a_{\tilde{n}}^{\tilde{m}},s_{\tilde{n}+1}^{\tilde{m}})$, but $r_{n+1}^{m}\neq r_{\tilde{n}+1}^{\tilde{m}}$.111 While we do not consider stochastic or noisy rewards in this work, we posit reward conflicts might be detected by sampling histories, in these cases.
-For example, if we define the history of $\tau_{m,n}$ to be $h(\tau_{m,n})=\{(s_{1}^{m},a_{1}^{m},r_{2}^{m},s_{2}^{m}),(s_{2}^{m},a_{2}^{m},r_{3}^{m},s_{3}^{m}),...,(s_{n-1}^{m},a_{n-1}^{m},r_{n}^{m},s_{n}^{m})\}$, then if $\tau_{m,n}$ is observed to conflict with $\tau_{\tilde{m},\tilde{n}}$ we might continually sample over $h(\tau_{m,n})$ and $h(\tau_{\tilde{m},\tilde{n}})$ to assess if their reward distributions are different in some statistical sense (e.g., difference of means).
+We define a reward conflict to be any two $\tau_{m,n}$ and $\tau_{\tilde{m},\tilde{n}}$ where $(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=(s_{\tilde{n}}^{\tilde{m}},a_{\tilde{n}}^{\tilde{m}},s_{\tilde{n}+1}^{\tilde{m}})$, but $r_{n+1}^{m}\neq r_{\tilde{n}+1}^{\tilde{m}}$.
 The motivation of our ILP is to map $T_{o}$ onto a valid ARMDP such that we resolve all reward conflicts observed in $T_{o}$.
-Intuitively, $\tilde{S}$, provides a degree of freedom over $U$, to do just that.
+Intuitively, $\tilde{S}$ provides a degree of freedom over $U$ to do just that.
 We define the matrix, $O=M\times N\times K\times K$, where $M=|T_{o}|$, $N$ reflects the largest $l_{m}$, $\forall\tau_{m}\in T_{o}$, and $K=|U|$ reflects the size of the finite state space for some candidate RM.
@@
 Each $O_{m,n,i,j}\in O_{m,n}$ is a binary variable reflecting the decision to map $s_{n}^{m}$ to RM state $u_{i}\in U$ and $s_{n+1}^{m}$ to RM state $u_{j}\in U$.
-That is, if $O_{m,n,i,j}=1$, then $\tau_{m,n}=(s_{n}^{m},a_{n}^{m},r_{n+1}^{m},s_{n+1}^{m})$ is mapped onto the ARMDP tuple, $((s_{n}^{m},u_{i}),a_{n}^{m},r_{n+1}^{m}(s_{n+1}^{m},u_{j}))$.
+That is, if $O_{m,n,i,j}=1$, then $\tau_{m,n}=(s_{n}^{m},a_{n}^{m},r_{n+1}^{m},s_{n+1}^{m})$ is mapped onto the corresponding ARMDP tuple of the form, $((s_{n}^{m},u_{i}),a_{n}^{m},r_{n+1}^{m}(s_{n+1}^{m},u_{j}))$.
+As such, each $\tau_{m,n}$ introduces $K^{2}$ variables through its respective $O_{m,n}$ matrix.
 We constrain each $O_{m,n}$ to choose a single mapping by enforcing that $\forall m=1,2,...,M,\forall n=1,2,...,l_{m}$:
 $\sum_{i=1}^{K}\sum_{j=1}^{K}O_{m,n,i,j}=1$ (6)
-We further constrain $O_{m,n}$ in instances where $n=1$ to satisfy the initial state requirement of Property 2 in the ARMDP. That is, $\forall m=1,2,...,M$:
+We further constrain $O_{m,n}$ where $n=1$ to satisfy the initial state requirement of Property 2 in the ARMDP. That is, $\forall m=1,2,...,M$:
 $\sum_{j=1}^{K}O_{m,1,1,j}=1$ (7)
 Next, we ensure the continuity of each trajectory in the mapping process.
-That is, we encode agreement between subsequent mappings of $O_{m,n}$ and $O_{m,n+1}$.
-To this end, we enforce the following constraint $\forall m=1,2,...,M,\forall n=1,2,...,l_{m}-1,\forall j=1,2,...,K$:
-$\sum_{i=1}^{K}O_{m,n,i,j}=\sum_{i=1}^{K}O_{m,n+1,j,i}$ (8)
-Note that we swap indices $j$ and $i$ on the right-hand side.
-Equation (8) simply forces $s_{n+1}^{m}$ to be mapped to the same RM state in both $\tau_{m,n}$ and $\tau_{m,n+1}$.
-Eqs (6-8) form the basic scaffolding of our mapping.
+That is, we encode agreement between subsequent mappings of all $O_{m,n}$ and $O_{m,n+1}$ pairs by enforcing the following constraint $\forall m=1,2,...,M,\forall n=1,2,...,l_{m}-1,\forall j=1,2,...,K$:
+$\sum_{i=1}^{K}O_{m,n,i,j}=\sum_{j^{\prime}=1}^{K}O_{m,n+1,j,j^{\prime}}$ (8)
+Equation (8) simply forces $s_{n+1}^{m}$ to be mapped to the same RM state in both $\tau_{m,n}$ and $\tau_{m,n+1}$. Equations (6-8) form the scaffolding of our mapping.
 Thus far, all constraints have been applied over individual or subsequent tuples of $T_{o}$.
@@
 $I_{s,a,s^{\prime},i,j}=\left[\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1,\\ (s,a,s^{\prime})=(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\end{subarray}}^{l_{m}}O_{m,n,i,j}\geq 1\right]$
-$I_{s,a,s^{\prime},i,j}$ is an indicator variable over the sum of all $O_{m,n,i,j}$ in $O$ with transition tuple, $(s,a,s^{\prime})$, and assignments $i,j\in\{1,2,...,K\}$.
-If $I_{s,a,s^{\prime},i,j}=1$, then our resulting ARMDP encodes the transition $((s,u_{i}),a,(s^{\prime},u_{j}))$ in $\tilde{T}$, and thus, $\delta_{u}(u_{j},L(s,a,s^{\prime}))=u_{j}$.
-If we let $(S,A,S)_{o}=\{(s_{n}^{m},a_{n}^{m},s_{n+1}^{m}),\forall\tau_{m,n}\in\tau_{m},\forall\tau_{m}\in T_{o}\}$ be the set of all unique transitions in $T_{o}$, then to enforce the determinism of $\delta_{u}$ we impose that $\forall(s,a,s^{\prime})\in(S,A,S)_{o},\forall i=1,2,...,K$:
+$I_{s,a,s^{\prime},i,j}$ is an indicator variable representing the sum of all $O_{m,n,i,j}$ in $O$ sharing the transition tuple $(s,a,s^{\prime})$ with assignments $i,j\in{1,2,...,K}$.
+When $I_{s,a,s^{\prime},i,j}=1$, it acts as a hidden trigger that highlights the sensitivity of $L$ to $(s,a,s^{\prime})$ and enforces the transition $\delta_{u}(u_{j},L(s,a,s^{\prime}))=u_{j}$, in the RM, thus governing $\tilde{T}$.
+We later consider the sensitivity of hidden triggers to rewards in Property 4.
+Let $(S,A,S)_{o}=\{(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\mid m\in\{1,2,...,M\},n\in\{1,2,...,l_{m}\}\}$ be the set of all unique transitions in $T_{o}$.
+Then to enforce the determinism of $\delta_{u}$, we impose that $\forall(s,a,s^{\prime})\in(S,A,S)_{o},\forall i=1,2,...,K$:
 $\sum_{j=1}^{K}I_{s,a,s^{\prime},i,j}\leq 1$ (9)
-Consider that for some $(s,a,s^{\prime})\in(S,A,S)_{o}$, and for some fixed $i\in\{1,2,...,K\}$ that when we constrain the sum $\forall j=1,2,...,K$ to be less than or equal to 1, we limit $\delta_{u}(u_{i},L(s,a,s^{\prime}))$ to have at most one outcome.
-We note that Equation (9) would be an equality when spanning $S\times A\times S$ instead of $(S,A,S)_{o}$.
-Equation (9), then, satisfies the determinism of $\delta_{u}$ and, therefore, Property 3 of the ARMDP.
+Consider that for any $(s,a,s^{\prime})\in(S,A,S)_{o}$ and for any $i\in\{1,2,...,K\}$ that when we constrain the sum $\forall j=1,2,...,K$, in Equation (9), to be less than or equal to 1, we limit $\delta_{u}(u_{i},L(s,a,s^{\prime}))$ to have at most one outcome.
+With this, Equation (9) satisfies the determinism of $\delta_{u}$ and Property 3 of the ARMDP.
 The encoding of Property 4 enforces the resolution of reward conflicts in $T_{o}$.
@@
 $I_{s,a,r,s^{\prime},i,j}=\left[\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1,\\ (s,a,r,s^{\prime})=\tau_{m,n}\end{subarray}}^{l_{m}}O_{m,n,i,j}\geq 1\right]$
-$I_{s,a,r,s^{\prime},i,j}$ is an indicator variable over the sum of all $O_{m,n,i,j}$ in $O$ with reward tuple, $(s,a,r,s^{\prime})$, and assignments $i,j\in\{1,2,...,K\}$.
-If $I_{s,a,r,s^{\prime},i,j}=1$, the ARMDP encodes $\tilde{R}((s,u_{i}),a,(s^{\prime},u_{j}))=r$, which is equivalent to $\delta_{r}(u_{i},L(s,a,s^{\prime}),u_{j})=r$ in the RM.
-If we let $r(s,a,s^{\prime})=\{r_{n}^{m},\forall\tau_{m,n}\in\tau_{m},\forall\tau_{m}\in T_{o}\mid(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=(s,a,s^{\prime})\}$ be the set of all observed rewards for some $(s,a,s^{\prime})$, then to enforce the determinism of $\delta_{r}$ we impose that $\forall(s,a,s^{\prime})\in(S,A,S)_{o},\forall i=1,2,...,K,\forall j=1,2,...,K$:
+$I_{s,a,r,s^{\prime},i,j}$ is an indicator variable representing the sum over all $O_{m,n,i,j}$ in $O$ sharing the reward tuple $(s,a,r,s^{\prime})$ with assignments $i,j\in\{1,2,...,K\}$.
+When $I_{s,a,r,s^{\prime},i,j}=1$, it imposes the following reward emission, $\delta_{r}(u_{i},L(s,a,s^{\prime}),u_{j})=r$, in the RM.
+Let $r(s,a,s^{\prime})=\{r_{n+1}^{m}\mid m\in\{1,2,...,M\},\hskip 2.0ptn\in\{1,2,...,l_{m}\},\\ (s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=(s,a,s^{\prime})\}$ be the set of all observed rewards for some $(s,a,s^{\prime})$.
+Then to enforce the determinism of $\delta_{r}$, we impose that $\forall(s,a,s^{\prime})\in(S,A,S)_{o},\forall i=1,2,...,K,\forall j=1,2,...,K$:
 $\sum_{r\in r(s,a,s^{\prime})}I_{s,a,r,s^{\prime},i,j}\leq 1$ (10)
-Consider that for some $(s,a,s^{\prime})\in(S,A,S)_{o}$ and for some fixed $i,j\in\{1,2,...,K\}$ that when we constrain the sum $\forall r\in r(s,a,s^{\prime})$ to be less than or equal to 1, we limit $\delta_{r}(u_{i},L(s,a,s^{\prime}),u_{j})$ to have at most one outcome, thus resolving our reward conflicts.
-Importantly, by Equation (10), we encode rewards as conditionals for inferring the structure of the RM.
-Therefore, if upstream rewards form dependencies with downstream rewards, rewards serve as memory, hence $H=(S\times A\times\mathbb{R})$.
-Like in Equation (9), Equation (10) would be an equality when spanning $\mathbb{R}$ and $S\times A\times S$ instead of some finite observed space.
-Regardless, by Equation (10) we satisfy the determinism of $\delta_{r}$ and, thereby, Property 4 of the ARMDP.
-Eqs (6-10) form the entirety of the constraints for ARMDP learning.
-However, there are potentially many ARMDP solutions for some given $T_{o}$.
+Consider that for any $(s,a,s^{\prime})\in(S,A,S)_{o}$ and for any $i,j\in\{1,2,...,K\}$ that when we constrain the sum $\forall r\in r(s,a,s^{\prime})$, in Equation (10), to be less than or equal to 1, we limit $\delta_{r}(u_{i},L(s,a,s^{\prime}),u_{j})$ to have at most one reward outcome, thus resolving our reward conflicts in $T_{o}$.
+Importantly, $I_{s,a,r,s^{\prime},i,j}=1$ contextualizes its corresponding hidden trigger, $I_{s,a,s^{\prime},i,j}$, with reward.
+When rewards are interdependent (e.g., some upstream reward conflict determines a downstream reward conflict), rewards serve as memory, hence, we capture relationships in $H=(S\times A\times\mathbb{R})^{*}$.
+By Equation (10), we satisfy the determinism of $\delta_{r}$ and Property 4 of the ARMDP.
+Equations (6-10) form the entirety of the constraints for ARMDP learning.
+However, there are potentially many ARMDP solutions for any given $T_{o}$.
 As such, we are motivated to derive simple and interpretable models for representing reward dynamics.
-Therefore, we orient our objective and learning procedure towards finding the most minimal ARMDP, and thus the most minimal RM, that satisfies $T_{o}$.
+Therefore, we orient our objective and learning procedure toward finding the minimal ARMDP, and thus the minimal RM, that satisfies $T_{o}$.
 This is done in two ways.
@@
 This is known as a deepening strategy employed in various related works [9, 6].
-Second, we limit the number of transition tuples that trigger transitions between RM states (e.g., $\delta_{u}(u_{i},L(s,a,s^{\prime}))=u_{j}$ where $u_{i}\neq u_{j}$).
+Second, we limit the number of hidden triggers that dictate transitions between different RM states (i.e., $\delta_{u}(u_{i},L(s,a,s^{\prime}))=u_{j}$ where $u_{i}\neq u_{j}$).
 We do so by minimizing the following objective:
 $z=\sum_{(s,a,s)\in(S,A,S)_{o}}\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1,\\ i\neq j\end{subarray}}^{K}I_{s,a,s^{\prime},i,j}$ (11)
-We solve Eqs (6-11) using out-of-the-box ILP solvers in Gurobi [14] to infer viable ARMDPs.
+We solve Equations (6-11) using out-of-the-box ILP solvers in Gurobi [13] to infer viable ARMDPs.
 We note that all notation in this section was constructed assuming $L:S\times A\times S\rightarrow 2^{P}$ though it is straightforward to rewrite it as being sensitive to $L:S\rightarrow 2^{P}$.
@@
 The implication of this is that we assume the RM to never return to a previous state, once left.
-While this might seem to limit the expressiveness of any resulting ARMDP we note that our ILP simply unrolls these cycles, under this assumption, and expands $U$ as needed.
-Because $T_{o}$ represents a finite set of finite trajectories, learned ARMDPs will be equivalently expressive, but larger in some cases.
-However, for all RMs considered in this work (e.g., the RMs in Figure 1) this structural assumption holds and we still learn the minimal ARMDP while dramatically reducing learning time.
+While this might seem to limit the expressiveness of any resulting ARMDP, we note that our ILP simply unrolls these cycles expanding $U$ as needed.
+Because $T_{o}$ represents a finite set of finite length trajectories, learned ARMDPs will be equivalently expressive, but larger in some cases.
+For all RMs considered in this work, this structural assumption holds resulting in the minimal ARMDP being learned while dramatically reducing learning time.
 We provide more details on this assumption in Appendix B.
@@
 Given that $K=|U|$ is determined by our deepening strategy, we assume $U$ is given.
-Additionally, by Equation (7) we designate $u_{1}\in U$, to be the initial state.
+Additionally, by Equation (7) we designate $u_{1}\in U$ to be the initial state.
 To extract the rules of $\delta_{u}$ and $\delta_{r}$ we simply iterate $\forall m=1,2,...M,\forall n=1,2,...,l_{m},\forall i=1,2,...,K,\forall j=1,2,...,K$.
@@
 4.4 Theoretical Considerations
-Our ILP is constructed such that it infers an ARMDP that satisfies reward conflicts in $T_{o}$.
-However, because the ARMDP is simply just an MDP, we can apply out-of-the-box RL algorithms to solve it.
-However, it remains to be shown how the ARMDP might be used as a proxy of the NMRDP for the purpose of maximizing reward expectations.
-Because we assume the form of $R$ in the NMRDP to be an RM, then any observed reward in $T_{o}$ would be consistently reproducible assuming identical histories.
-Therefore, so long as the likelihood of producing trajectories under $T$ is identical to $\tilde{T}$, reward expectations are preserved.
-We start with the weighted reward sum for some $\tau_{m}\in T_{o}$:
-$R(\tau_{m})=\sum_{n=1}^{l_{m}}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})$ (12)
-Similarly, we construct a weighted reward sum for some $\tilde{\tau}_{m}$ derived by mapping $\tau_{m}$ onto an ARMDP to be:
-$R(\tilde{\tau}_{m})=\sum_{n=1}^{l_{m}}\sum_{i=1}^{K}\sum_{j=1}^{K}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\cdot O_{m,n,i,j}$ (13)
-We remind readers that every $O_{m,n,i,j}$ reflects a binary variable that, if 1, allows us to conclude $\delta_{u}(u_{i},L(s_{n}^{m},a_{n}^{m},s_{n+1}^{m}))=u_{j}$.
-By referencing Property 3 of the ARMDP, the interpretation of multiplying $O_{m,n,i,j}$ into $T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})$ to determine $\tilde{T}$ should be evident.
-With this we introduce Theorem 1:
+Our ILP is constructed such that it infers an ARMDP that resolves reward conflicts in $T_{o}$.
+Because the ARMDP is simply just an MDP, we can apply out-of-the-box RL algorithms to solve it.
+However, it remains to be shown how the ARMDP might be a suitable proxy for the NMRDP for the purpose of maximizing reward expectations.
+We start with the weighted Reward Sum (RS) for some $\tau_{m}\in T_{o}$:
+$RS(\tau_{m})=\sum_{n=1}^{l_{m}}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})$ (12)
+Given that we assume the form of $R$ to be an RM and that $\delta_{u}$ and $\delta_{r}$, in an RM, are deterministic functions, $RS(\tau_{m})$ reliably captures the reward observations essential for constructing a reward expectation, assuming the trajectory sequence is preserved.
+Similarly, we construct a weighted Abstracted Reward Sum (ARS) for some $\tau_{m}\in T_{o}$ under the representation of the ARMDP to be:
+$ARS(\tau_{m})=\sum_{n=1}^{l_{m}}\sum_{i=1}^{K}\sum_{j=1}^{K}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\cdot O_{m,n,i,j}$ (13)
+We remind readers that every $O_{m,n,i,j}$ reflects a binary variable that, if 1, indicates $\delta_{u}(u_{i},L(s_{n}^{m},a_{n}^{m},s_{n+1}^{m}))=u_{j}$.
+By referencing Property 3 of the ARMDP, multiplying $O_{m,n,i,j}$ into $T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})$ to determine $\tilde{T}$ should be evident.
+We introduce Theorem 1:
 Theorem 1.
-$R(\tau_{m})=R(\tilde{\tau}_{m})$ $\forall\tau_{m}\in T_{o}$ where $\tilde{\tau}_{m}$ is derived by mapping $\tau_{m}$ onto an ARMDP.
-For space reasons, we provide the proof of Theorem 1 in Appendix C, but state that by Theorem 1, the ARMDP is a suitable proxy for maximizing reward sums $\forall\tau_{m}\in T_{o}$.
+$RS(\tau_{m})=ARS(\tau_{m})$ $\forall\tau_{m}\in T_{o}$.
+For space reasons, we provide the proof of Theorem 1 in Appendix C, but briefly state that by ensuring $RS(\tau_{m})=ARS(\tau_{m})$ $\forall\tau_{m}\in T_{o}$, the ARMDP is a suitable proxy for maximizing reward expectations for the NMRDP, assuming $T_{o}$ is representative.
+Furthermore, even if $T_{o}$ is not representative, we posit that any ARMDP derived from $T_{o}$ serves as a useful hypothesis for RL agents to interrogate in an active learning paradigm.
+We explore that intuition in the next section.
 4.5 RL for Non-Markov Rewards
 Our ILP formulation for learning ARMDPs is a form of passive learning.
-That is to say, the ARMDP will be sufficient for $T_{o}$, but may not reflect the ground truth.
-However, by Theorem 1, we know that we can use ARMDPs as a proxy for maximizing reward expectations for the NMRDP.
-Therefore, we apply Q-Learning to make our approach an active learner.
-Our intuition is that we might resolve the reward conflicts that are needed to maximize our expected reward.
+That is to say, while the ARMDP will sufficiently resolve reward conflicts found in $T_{o}$, $T_{o}$ may not be representative of the true reward dynamics.
+Therefore, we are interested in extending our approach into an active learning framework by applying Q-Learning over inferred ARMDPs.
+Our intuition is that while any instance of $T_{o}$, or any ARMDP inferred from it, may not be truly representative, Q-Learning agents might gradually expand $T_{o}$ in their pursuit to maximize reward expectations.
+In this way, each inferred ARMDP acts as a hypothesis that can be interrogated for consistency.
 Algorithm 1 ARMDPQ-Learning
 Input: NMRDP
-1:  $ARMDP=ARMDP_{0}$
-2:  $\pi_{ARMDP}=\pi_{0}$
+[…]
+[…]
 3:  $T_{o}=[]$
@@
 5:  while $!done$ do
-6:     $\tau_{m},conflict=NMRDP.sim(ARMDP,\pi_{ARMDP})$
-7:     if $conflict$ then
+[…]
+[…]
 8:        $T_{o}.append(\tau_{m})$
-9:        $solved=False$
-10:        while $!solved$ do
-11:           $ARMDP,solved=solve\_ILP(T_{o},K)$
+[…]
+[…]
+[…]
 12:           if $!solved$ then
@@
 15:        end while
-16:        $\pi_{ARMDP}=\pi_{0}$
+[…]
 17:     else
-18:        $\pi_{ARMDP}=update\_qsa(\tau_{m})$
+[…]
 19:     end if
 20:  end while
-We present ARMDPQ-Learning, with pseudocode available in Algorithm 1.
-The agent applies our deepening strategy by initially assuming the problem is an MDP (e.g., $|U|=1$).
-We denote this ARMDP as ARMDP0 (line 1).
-$K=|U|=2$, is used for subsequent ARMDP inference.
-The agent policy, $\pi_{ARMDP}$, is set to $\pi_{0}$, a policy with all Q-values set to 0 (line 2).
-We use the NMRDP to simulate some $\tau_{m}$ under the perspective of ARMDP, selecting actions according to $\pi_{ARMDP}$ (line 6).
-If a conflict is found we add $\tau_{m}$ into $T_{o}$ (line 8) and solve for a new ARMDP (line 11).
-If the problem is infeasible we increment $K$ by 1 (line 13) until there is a solution.
-As each ARMDP presents a new perspective, we reset $\pi_{ARMDP}$ back to $\pi_{0}$ (line 16).
-Assuming no conflict is found for $\tau_{m}$, we update $\pi_{ARMDP}$ with $\tau_{m}$ according to Q-Learning updating (line 18).
-Note that line 18 is shorthand.
-In reality, tuples are updated in an online fashion while simulating the trajectory.
-Also, by only adding $\tau_{m}$ known to conflict with our current ARMDP, we keep $T_{o}$ small to make our ILP solving more efficient.
+[…]
 5 Experiments
@@
 5.1 Experiment 1 - Officeworld
-We validate our approach by applying ARMDPQ-Learning on the four black-box RMs (Tasks (b-e)) from the Officeworld domain visualized in Figure 1.
-We extend each RM with a terminal state on all * tiles, resulting in a Markov reward of -1.
-The RM, as well as $P$ and $L$ are hidden from the agent.
-We refer readers to Appendix D for details on hyper-parameter selection, but note that we use an $\epsilon$-greedy policy and set $\epsilon$ to 0% after 90,000 episodes to exploit learned behavior.
+We validated our approach by applying ARMDPQ-Learning on the four black-box RMs (Tasks (b-e)) from the Officeworld domain visualized in Figure 1.
+We hid the RM as well as $P$ and $L$ from the agent.
+We refer readers to Appendix D for details on hyper-parameter selection, but note that we used an $\epsilon$-greedy policy and set $\epsilon$ to 0% after 90,000 episodes to exploit learned behavior.
 Results for Experiment 1 can be seen in Figure 2.
-Figure 2: ARMDPQ-Learning evaluated on Tasks (b-e) in the Officeworld domain on 10 random trials. Rewards show +/- a standard deviation.
-The agent learns to perform optimally in all tasks, suggesting that a meaningful representation was learned.
+Figure 2: Average rewards with standard deviations, evaluated over 10 trials using ARMDPQ-Learning on Tasks (b-e) in the Officeworld domain.
+The agent learned to perform optimally in all tasks, suggesting that a meaningful representation was learned.
 Task (b), (c), and (e) all arrived at $|U|$ reflecting the ground truth model (2, 2, and 4, respectively).
@@
 $T_{o}$ remained relatively small for all tasks, with $|T_{o}|\approx$ 4, 8, 39, and 44 for Tasks (b-e), respectively.
-We also tracked the cumulative solve times for each task to be $\approx$ 0.1, 0.6, 4,200, and 12,000 seconds, respectively.
+We also tracked the cumulative solve times for all ILP solves under each task to be $\approx$ 0.1, 0.6, 4,200, and 12,000 seconds, respectively.
 We include an extension to Figure 2 in Appendix D, containing figures for $|U|$, $|T_{o}|$ and cumulative solve times.
@@
 We attribute this to identifying the wrong pattern in $H$ during early conflict resolution.
-Interestingly, because these early solutions are so poor, they are easily testable by the Q-Learing agent.
+Interestingly, because these early solutions were so poor, they were easily testable by the Q-Learning agent.
 This lends some credence to our initial intuition of the Q-Learning agent serving as an important active learner capable of interrogating its own perspectives and updating in the face of new evidence.
@@
 Previously, Christoffersen et al.
-compared the learning rates of Q-Learning agents over NMRDPs, using DFAs, to sparse, recurrent deep RL models [9].
+compared the learning rates of Q-Learning agents over NMRDPs, using DFAs, to sparse, recurrent, deep RL models [9].
 However, in smaller discrete state spaces, Q-Learning has the advantage of updating state-action values directly given their simpler, tabular representation.
 In contrast, deep RL models must learn approximate solutions using complex computations over large batches of data.
-It remains to be shown how beneficial it could be to represent the state space of a DQN with the state space of an ARMDP.
-This would allow for a more even comparison in learning performance against RDQN architectures that have to learn a Markov representation over the sparse histories they encode.
-For each Task (b-e), we extract the ARMDP representations from Experiment 1 and represent $S$ and $U$ as the one-hot encoded feature space of an Abstracted DQN (ADQN).
-As the ADQN explores each task we update $S$ and $U$ according to its respective ARMDP.
-In parallel, we train recurrency-based models without access to these abstractions, namely, a Recurrent DQN (RDQN), a Long-Short Term Memory DQN (LSTMDQN), and a Gated Recurrent DQN (GRUDQN).
-For a baseline, we also train a DQN without any knowledge of $U$.
-Figure 3: DQN models evaluated on Tasks (b-e) from the Officeworld domain on 10 random trials.
-Rewards are very sparse for the Officeworld tasks and exploration via an $\epsilon$-greedy policy can often be insufficient.
-As we don’t seek to demonstrate the efficacy of RL strategies under sparse rewards, we occasionally seed each model with an optimal trajectory (every 10th trajectory).
+It remained to be shown how beneficial it could be to represent the state space of a DQN with $\tilde{S}$ instead of $S$.
+This would allow for a more even comparison of training profiles against RDQN architectures that have to learn a Markov representation over histories of $S$.
+For each Task (b-e), in Experiment 1, we extracted $\tilde{S}$, from the finalized ARMDP.
+We used $\tilde{S}$ as the feature space for an Abstracted DQN (ADQN).
+As the ADQN explored each task we updated $\tilde{S}$ according to its respective ARMDP.
+In parallel, we trained recurrency-based models using $S$, namely, an RDQN, a Long-Short Term Memory DQN (LSTMDQN), and a Gated Recurrent DQN (GRUDQN).
+For a baseline, we also trained a DQN using $S$.
+Figure 3: Average rewards with standard deviations, evaluated over 10 trials using DQN models on Tasks (b-e) in the Officeworld domain.
+Rewards were very sparse for the Officeworld tasks and exploration via an $\epsilon$-greedy policy was often insufficient.
+As we didn’t seek to demonstrate the efficacy of RL strategies under sparse rewards, we seeded each model with an optimal trajectory every 10th trajectory.
 Results for Experiment 2 can be seen in Figure 3.
-In all tasks, the ADQN model learns at the rate of, or out-learns, all other models.
-This becomes more evident as the complexity of the task increases (Task (e) being the most complex).
-We note that for Task (b), the baseline DQN performs just as well as the other models only because the shortest path to the goal includes the coffee tile.
-Interestingly, for Task (e), the RDQN performs just as poorly as the baseline DQN model while the LSTMDQN and GRUDQN models (while learning slowly) appear to be uncovering the relevant hidden mechanisms of the underlying RM.
+In all tasks, the ADQN model learned at the rate of, or out-learned, all other models.
+This became more apparent as the complexity of the task increased (Task (e) being the most complex).
+We note that for Task (b), the baseline DQN performed just as well as the other models only because the shortest path to the goal included the coffee tile.
+Interestingly, for Task (e), the RDQN performed just as poorly as the baseline DQN model while the LSTMDQN and GRUDQN models (while learning slowly) appeared to be uncovering the relevant hidden mechanisms of the underlying RM.
 We hypothesize that the forgetting component of the GRUDQN and LSTDQN is a key differentiating factor in their learning ability.
-Importantly, we remind readers to still consider the cumulative solve times of ARMDQ-Learning from Experiment 1 as this is a non-trivial time cost.
+Still, we remind readers that uncovering $U$, as discussed in Experiment 1, involves a non-trivial cost before ADQN training can commence.
 5.3 Experiment 3 - Breakfastworld
@@
 $(u_{1},c,u_{2})$ returns a reward of -0.1, reflecting a time cost.
-However, this time cost is experienced in a non-linear fashion as the agent grows bored and spending another turn cooking returns a reward of -0.3, $(u_{2},c,u_{3})$.
-However, this patience pays off in $(u_{3},e,u_{4})$, where eating a perfectly cooked meal returns a reward of 2.
+This time cost is experienced in a non-linear fashion as the agent grows bored and spending another turn cooking returns a reward of -0.3, $(u_{2},c,u_{3})$.
+However, this patience pays off in $(u_{3},e,u_{4})$, where eating a well cooked meal returns a reward of 2.
 In contrast, $(u_{2},e,u_{5})$ returns 1, reflecting an undercooked meal.
-Finally, $(u_{4},l,term)$ and $(u_{5},l,term)$ reflect the last reward dependency, returning the cumulative reward for each path in the RM.
+Finally, $(u_{4},l,term)$ and $(u_{5},l,term)$ reflect the last reward dependency, returning the cumulative reward for each path in the RM, encouraging episode completion.
 Task (c) builds on Task (b) by adding a reward dependency for washing.
@@
 Both tasks reflect RMs more complex in terms of $|U|$ than Task (e) in Experiment 1.
-However, we will demonstrate that RM learning benefits in cases where rewards are interdependent, as observed reward conflicts result in natural splits in the ARMDP learning by Equation (10).
-Conversely, DFA learning suffers under such conditions as each reward conflict must be represented by its own DFA [13].
-Moreover, downstream DFAs are learned independently from upstream DFAs as they only learn according to $H=(S\times A)^{*}$.
+However, we demonstrate that RM learning benefits in cases where rewards are interdependent as upstream reward conflict resolution results in natural splits for downstream reward conflict resolution.
+Conversely, DFA learning suffers under such conditions as each reward conflict must be represented by its own DFA which is learned independently from any other DFA [12], hence they only learn according to $H=(S\times A)^{*}$.
 We demonstrate how advantageous it can be that our ILP naturally learns according to $H=(S\times A\times\mathbb{R})^{*}$ by applying ARMDPQ-Learning to Tasks (b) and (c).
-We then compare this to applying ARMDPQ-Learning on RMs equivalent in structure to Task (b) and (c), but that only emit the cumulative rewards on the terminal state.
+We then compare this to applying ARMDPQ-Learning on RMs equivalent in structure to Task (b) and (c), but that only emit cumulative rewards on terminal transitions.
 All other rewards are made 0.
 We refer to Tasks (b) and (c), in Figure 4, as Task (b) Full and Task (c) Full, respectively.
-Conversely, the RMs that only emit the cumulative reward are referred to as Task (b) Cumulative and Task (c) Cumulative, respectively.
-The purpose of these cumulative tasks is to form a problem that produces reward conflicts on tile [0,0] that can not be contextualized by the prior reward conflicts.
-This would be similar to learning two DFAs to parse the two different possible rewards on tile [0,0] only using $H=(S\times A)^{*}$.
+Conversely, the RMs that only emit the cumulative rewards are referred to as Task (b) Cumulative and Task (c) Cumulative, respectively.
+The purpose of these cumulative tasks is to form a problem that produces reward conflicts only on the $l$ tile that can not be contextualized by upstream reward conflicts.
+This would be similar to learning DFAs to parse the different possible rewards on the $l$ tile only using $H=(S\times A)^{*}$.
 We show results for Experiment 3 in Table 1.
-In all solvable cases of ARMDP-Learning, the agent achieved optimal performance and identified the correct $|U|$ for each task.
+In all solvable cases of ARMDPQ-Learning, the agent achieved optimal performance and identified the correct $|U|$, for each task.
 We include an extension figure to Table 1 in Appendix F containing such results.
-Using Task (b) Full our cumulative solve time ($\approx$ 4.4 seconds) was dramatically faster than when using Task (b) Cumulative ($\approx$ 1,102.9 seconds).
-This can be attributed to Task (b) Full requiring roughly a fourth of the number of trajectories to solve than Task (b) Cumulative, likely because of the leveraged context of upstream rewards.
+Using Task (b) Full our cumulative solve time for all ILPs was dramatically smaller ($\approx$ 4.4 seconds) than when using Task (b) Cumulative ($\approx$ 1,102.9 seconds).
+This can be attributed to Task (b) Full requiring only a fourth of the number of trajectories to solve than Task (b) Cumulative, likely due to the leveraged context of upstream rewards.
 More impressive were the results for Task (c) Full which resulted in the cumulative solve time of $\approx$ 153.8 seconds.
@@
 The ARMDP represents the cross-product between an RM and the underlying state space of an NMRDP.
-Importantly, we learn the ARMDP, and thus the RM, without knowledge of $P$, by identifying hidden triggers that govern the rules of $L(S\times A\times S)\rightarrow 2^{P}$.
-We also demonstrated how effective it can be to learn our automaton according to $H=(S\times A\times\mathbb{R})^{*}$, rather than $H=(S\times A)^{*}$, a limitation of DFAs.
-We proved that the ARMDP shares an equivalent expectation to the NMRDP for $T_{o}$ and can therefore be used as a proxy for maximizing reward expectations.
+Importantly, we learn the ARMDP, and thus the RM, without knowledge of $P$ or $L$, by identifying hidden triggers that govern the rules of $L:(S\times A\times S)\rightarrow 2^{P}$.
+We also demonstrated how effective it can be to learn our automata according to $H=(S\times A\times\mathbb{R})^{*}$, rather than $H=(S\times A)^{*}$, a limitation of DFAs.
+We proved that the ARMDP is a suitable proxy for maximizing the reward expectations of the NMRDP for $T_{o}$.
 We validated our approach by applying ARMDPQ-Learning to train Q-Learning agents over black-box RMs.
@@
 We compared the training profiles of these DQNs to the training profiles of various RDQN architectures that had to learn Markov representations over the sparse histories they encode.
-One important aspect of this work is that the RMs we learn are inherently more interpretable than the sparse representation of histories in $H=(S\times A\times\mathbb{R})^{*}$.
-This is in part because we learn minimal RMs that encode $H$ under the finite, structured, and symbolic representation of an automaton.
-However, more importantly, interpreting an RM doesn’t require a comprehensive understanding of $S$ and $A$ outside of what the labeler encodes.
+An important aspect of this work is that the RMs we learn are inherently more interpretable than the sparse representation of histories in $H=(S\times A\times\mathbb{R})^{*}$.
+This is in part because we learn minimal RMs that encode $H$ under the finite, structured, and symbolic representation of an automaton, but also because interpreting an RM doesn’t require a comprehensive understanding of $S$ and $A$ outside of what the labeler encodes.
 This is a consequence of the RM being entirely decomposable from the ARMDP.
-For example, the RM in Figure 1 Task (b) conveys that we need to bring coffee to the office without ever needing to describe where the coffee might be sourced from or where the office is.
-Since RL is predicated on the assumption that the reward function is the most concise, robust, and transferable definition of a task [20], we posit that RMs emerge as strong candidates for reward function representation.
+For example, the RM in Figure 1 Task (b) conveys the need to bring coffee to the office without ever needing to describe the sequence of states and actions to achieve that goal.
+Since RL is predicated on the assumption that the reward function is the most concise, robust, and transferable definition of a task [19], we posit that RMs emerge as strong candidates for reward function representation and interpretation.
 RM representations also carry important weight for the field of Inverse RL (IRL).
-In IRL, the goal is to learn a reward function from behavioral data that motivated the behavior [20, 26, 8].
+In IRL, the goal is to learn a reward function from behavioral data that motivated that behavior [19, 25, 8].
 While most approaches assume a Markov reward function, some of this work has been focused on learning rewards in Partially Observable MDPs (POMDPs) [8] and in Semi MDPs (SMDPs) [25].
@@
 Lastly, we generally make note that this work is limited in that it assumes $S$ and $A$ are discrete.
-As such, we would like to consider automata extraction methods, in the future, using RNNs [28, 21, 31] as this would extend our approach into the continuous case.
+As such, we would like to consider automata extraction methods, in the future, using recurrent models [27, 20, 30] as this would extend our approach into the continuous case.
 Weiss et al.
-have demonstrated promising results in this regard by applying $L*$ as an exact learning algorithm for deep recurrent models [31].
+have demonstrated promising results in this regard [30].
+Similarly, this work assumes $R$ to be deterministic to some history.
+In scenarios where $R$ is stochastic or noisy, we posit that we might be able to handle such cases by sampling histories to assess if the reward distributions differ in some statistical sense.
+{ack}
+This research was funded in part by the Air Force Office of Scientific Research Grant No. FA9550-20-1-0032 and the Office of Naval Research Grant No. N00014-19-1-2211.
 Appendix
@@
 $\sum_{j=1}^{K}O_{m,1,1,j}=1$ $\forall m=1,2,...,M$ $\hskip 2.0pt(3)$
-$\sum_{i=1}^{K}O_{m,n,i,j}=\sum_{i=1}^{K}O_{m,n+1,j,i}$ $\forall m=1,2,...,M,\forall n=1,2,...,l_{m}-1,\forall j=1,2,...,K$ $\hskip 2.0pt(4)$
+$\sum_{i=1}^{K}O_{m,n,i,j}=\sum_{j^{\prime}=1}^{K}O_{m,n+1,j,j^{\prime}}$ $\forall m=1,2,...,M,\forall n=1,2,...,l_{m}-1,\forall j=1,2,...,K$ $\hskip 2.0pt(4)$
 $\sum_{j=1}^{K}I_{s,a,s^{\prime},i,j}\leq 1$ $\forall(s,a,s^{\prime})\in(S,A,S)_{o},\forall i=1,2,...,K$ $\hskip 2.0pt(5)$
@@
 $O_{m,n,ij}\in\{0,1\}$ ${\LARGE\begin{subarray}{c}\forall m=1,2,...,M,\forall n=1,2,...,l_{m},\\ \hskip 13.0pt\forall i=1,2,...,K,\forall j=1,2,...,K\end{subarray}}$ $\hskip 2.0pt(9)$
-The goal of our Integer Linear Program (ILP) formulation is to infer a valid Abstract Reward Markov Decision Process (ARMDP), as defined in Definition 1, from a set of observed trajectories with observed reward conflicts.
-Reward conflicts are a result of an agent observing rewards under the lens of $S\times A\times S$ when in reality rewards are sensitive to histories of the form $H=(S\times A\times\mathbb{R})^{*}$.
+The goal of our Integer Linear Program (ILP) formulation is to infer a valid Abstract Reward Markov Decision Process (ARMDP), as defined in Definition 1, in the main paper, from a set of observed trajectories, $T_{o}$, with observed reward conflicts.
+Reward conflicts are a result of an agent observing rewards under the lens of $S\times A\times S$, when in reality, rewards are sensitive to histories of the form $H=(S\times A\times\mathbb{R})^{*}$.
 We introduce some notation that our ILP leverages.
-We derive candidate ARMDPs by solving an ILP over a set of observed trajectory traces, $T_{o}$.
 Let, $\tau_{m}\in T_{o}$, be the m-th trace in $T_{o}$ of length $l_{m}$.
@@
 We define a reward conflict to be any two $\tau_{m,n}$ and $\tau_{\tilde{m},\tilde{n}}$ where $(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=(s_{\tilde{n}}^{\tilde{m}},a_{\tilde{n}}^{\tilde{m}},s_{\tilde{n}+1}^{\tilde{m}})$, but $r_{n+1}^{m}\neq r_{\tilde{n}+1}^{\tilde{m}}$.
-We define the matrix, $O=M\times N\times K\times K$, where $M=|T_{o}|$, $N$ reflects the largest $l_{m}$, $\forall\tau_{m}\in T_{o}$, and $K=|U|$ reflects the size of the finite state space for some candidate RM.
+We define the matrix, $O=M\times N\times K\times K$, where $M=|T_{o}|$, $N$ reflects the largest $l_{m}$, $\forall\tau_{m}\in T_{o}$, and $K=|U|$ reflects the size of the finite state space for some candidate Reward Machine (RM).
 $O_{m,n}$, then, is a $K\times K$ matrix that reflects the mapping of $\tau_{m,n}=(s_{n}^{m},a_{n}^{m},r_{n+1}^{m}s_{n+1}^{M})$ onto the ARMDP, where both $s_{n}^{m}$ and $s_{n+1}^{m}$ are assigned an RM state.
 Additionally, we make use of two sets to gather necessary variables from $T_{o}$ for the construction of our ILP.
-These are $(S,A,S)_{o}=\{(s_{n}^{m},a_{n}^{m},s_{n+1}^{m}),\forall\tau_{m,n}\in\tau_{m},\forall\tau_{m}\in T_{o}\}$ and $r(s,a,s^{\prime})=\{r_{n}^{m},\forall\tau_{m,n}\in\tau_{m},\forall\tau_{m}\in T_{o}\mid(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=(s,a,s^{\prime})\}$ where $(S,A,S)_{o}$ reflects the set of all observed transition tuples in $T_{o}$ and $r(s,a,s^{\prime})$ is the set of all observed rewards for some observed transition tuple in $T_{o}$.
+These are $(S,A,S)_{o}=\{(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\mid m\in\{1,2,...,M\},n\in\{1,2,...,l_{m}\}\}$ and $r(s,a,s^{\prime})=\{r_{n+1}^{m}\mid m\in\{1,2,...,M\},n\in\{1,2,...,l_{m}\},(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=(s,a,s^{\prime})\}$ where $(S,A,S)_{o}$ reflects the set of all observed transition tuples, $(s,a,s^{\prime})\in S\times A\times S$, that exist in $T_{o}$, and $r(s,a,s^{\prime})$ is the set of all observed rewards for some given $(s,a,s^{\prime})$.
 Appendix B j $\geq$ i Constraint
@@
 The implication of this construction on the form of inferred ARMDP is that the underlying RM can never return to a prior state, once left.
-To reiterate, this does not limit the expressivity of the inferred ARMDP, it simply requires larger $U$ in some instances.
+To reiterate our sentiments from the main paper, this does not limit the expressivity of the inferred ARMDP, it simply requires a larger $U$ in some instances.
 Consider the RMs in Figure 5.
-Task (a) represents a patrol task identical to Task (e) in Figure 1 in the main paper.
+Task (a) represents a patrol task identical to Task (e) in Figure 1, in the main paper.
 Even though Task (a) has atomic cycles (e.g., $u_{1}\rightarrow u_{1}$), once a state is left, the RM never returns to it.
-For RMs like this, even with $j\geq i$, we will still be able to represent the RM minimally.
-Conversely, Task (b) is a continuous patrol task, where on completion of a lap, the RM begins again in $u_{1}$.
+For RMs like this, even with $j\geq i$ imposed, we still learn the minimal RM.
+Conversely, Task (b) is a continuous patrol task, where on completion of a single patrol cycle, the RM begins again in $u_{1}$.
 Here, because the RM reuses previous states, the ARMDPs that we produce will continuously unroll these states and expand $U$ as needed.
@@
 Appendix C ARMDPs as NMRDP Proxies
-The goal of this section is to prove that the ARMDP is a sufficient proxy of the NMRDP for the purposes of maximizing reward expectations.
-Because we assume the form of $R$ in the NMRDP to be an RM, then any observed reward in $T_{o}$ would be consistently reproducible assuming identical histories.
-Therefore, so long as the likelihood of producing trajectories under $T$ is identical to $\tilde{T}$, reward expectations are preserved.
-We restate the the weighted reward sum for some $\tau_{m}\in T_{o}$:
-$R(\tau_{m})=\sum_{n=1}^{l_{m}}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})$ (14)
-Similarly, we restate the weighted reward sum for some $\tilde{\tau}_{m}$ derived by mapping $\tau_{m}$ onto an ARMDP to be:
-$R(\tilde{\tau}_{m})=\sum_{n=1}^{l_{m}}\sum_{i=1}^{K}\sum_{j=1}^{K}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\cdot O_{m,n,i,j}$ (15)
+The goal of this section is to further elaborate on how the ARMDP is a sufficient proxy of the NMRDP for the purposes of maximizing reward expectations.
+We start by reintroducing the weighted Reward Sum (RS) for some $\tau_{m}\in T_{o}$:
+$RS(\tau_{m})=\sum_{n=1}^{l_{m}}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})$ (14)
+Given that we assume the form of $R$ to be an RM and that $\delta_{u}$ and $\delta_{r}$, in an RM, are deterministic functions, $RS(\tau_{m})$ reliably captures the reward observations essential for constructing a reward expectation, assuming the trajectory sequence is preserved.
+Similarly, we construct a weighted Abstracted Reward Sum (ARS) for some $\tau_{m}\in T_{o}$ under the representation of the ARMDP to be:
+$ARS(\tau_{m})=\sum_{n=1}^{l_{m}}\sum_{i=1}^{K}\sum_{j=1}^{K}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\cdot O_{m,n,i,j}$ (15)
 We remind readers that every $O_{m,n,i,j}$ reflects a binary variable that, if 1, allows us to conclude $\delta_{u}(u_{i},L(s_{n}^{m},a_{n}^{m},s_{n+1}^{m}))=u_{j}$.
@@
 $\tilde{T}((s,u),a,(s^{\prime},u^{\prime}))=\begin{cases}T(s,a,s^{\prime})&\text{if\ }\delta_{u}(u,L(s,a,s^{\prime}))=u^{\prime}\\ 0&\text{otherwise}\end{cases}$
-By multiplying $O_{m,n,i,j}$ into $T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})$, $O_{m,n,i,j}$ serves as a logical gate to produce the dynamics of $\tilde{T}$. With this we introduce Theorem 1:
+By multiplying $O_{m,n,i,j}$ into $T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})$, $O_{m,n,i,j}$ serves as a logical gate to produce the dynamics of $\tilde{T}$. With this we reintroduce Theorem 1:
 Theorem 1.
-$R(\tau_{m})=R(\tilde{\tau}_{m})$ $\forall\tau_{m}\in T_{o}$ where $\tilde{\tau}_{m}$ is derived by mapping $\tau_{m}$ onto an ARMDP.
+$RS(\tau_{m})=ARS(\tau_{m})$ $\forall\tau_{m}\in T_{o}$.
 Proof.
-$R(\tilde{\tau}_{m})$ $=\sum_{n=1}^{l_{m}}\sum_{i=1}^{K}\sum_{j=1}^{K}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\cdot O_{m,n,i,j}$
-$=\sum_{n=1}^{l_{m}}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\sum_{i=1}^{K}\sum_{j=1}^{K}\cdot O_{m,n,i,j}$
+$ARS(\tau_{m})$ $=\sum_{n=1}^{l_{m}}\sum_{i=1}^{K}\sum_{j=1}^{K}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\cdot O_{m,n,i,j}$
+$=\sum_{n=1}^{l_{m}}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\sum_{i=1}^{K}\sum_{j=1}^{K}O_{m,n,i,j}$
 $=\sum_{n=1}^{l_{m}}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\cdot 1$
-$=\sum_{n=1}^{l_{m}}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=R(\tau_{m})$
+$=\sum_{n=1}^{l_{m}}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=RS(\tau_{m})$
+Assuming $T_{o}$ is representative sample of the reward dynamics of $R$, in an NMRDP, by Theorem 1, the ARMDP is a suitable proxy for maximizing reward expectations for the NMRDP.
 Appendix D Experiment 1 Specifications
 For all tasks in Experiment 1, we used $L:S\rightarrow 2^{P}$ and $j\geq i$ toggle construction.
-Hyper parameters include $\gamma=0.95$, $\alpha=0.1$ and $\epsilon=0.1$.
-The only exception is for Task (b) where we use $\epsilon=0.2$ until the first conflict, to promote exploring.
-This is because the shortest path to the goal state includes the necessary tile to trigger the non-Markov reward.
-As such, the Markov representation under $S$ and $A$ appears sufficient to demonstrate optimal behavior.
+We adapted the Officeworld environment from https://github.com/RodrigoToroIcarte/reward_machines.git.
+We noticed that in this implementation, the * tile was a terminal state but resulted in a neutral reward of 0.
+We instead marked all * tiles as terminal states with a reward of -1 to deter the agent from stepping onto them.
+Hyper parameters included $\gamma=0.95$, $\alpha=0.1$ and $\epsilon=0.1$.
+The only exception is for Task (b) where we used $\epsilon=0.2$ until the first conflict, to promote exploring.
+This is because the shortest path to the goal state included the necessary tile to trigger the non-Markov reward.
+As such, the Markov representation under $S$ and $A$ appeared sufficient to demonstrate optimal behavior.
 Even with $\epsilon=0.2$, for Task (b), 100,000 episodes was oftentimes not enough to trigger a conflict.
-To further accommodate this we put a sliding window on the total episode counts used to produce Figure 2 in the main paper such that the 100,000 episodes contained the first conflict.
+To further accommodate this we put a sliding window on the total episode counts used to produce Figure 2, in the main paper, such that the 100,000 episodes contained the first conflict.
 This was strictly to produce a consistent figure over 10 randomly seeded trials.
@@
 Machine specifications used in the experiment: 32 core Intel(R) Xeon(R) CPU E5-2640 v3 @ 2.60 GHz with 512 GB of Memory.
-Figure 6: Results of ARMDPQ-Learning evaluated on Tasks (b-e) from the Officeworld domain on 10 random trials.
-Episode Reward Avg column shows average reward +/- a stddev.
+Figure 6: Results of ARMDPQ-Learning evaluated over 10 random trials on Tasks (b-e) in the Officeworld domain.
+Average Rewards column is displayed with standard deviations.
 $|U|$ column shows current assumed size of $U$.
@@
 Appendix E Experiment 2 Specifications
-Each model in Experiment 2 was trained using a Double Deep Q-Network (DDQN) architecture, meaning each model had a target and non-target model supporting the Deep Q-Network (DQN) architecture.
+Each model in Experiment 2 was trained using a Double Deep Q-Network (DDQN) architecture, meaning that it leveraged a target and non-target model.
 All target models were trained slowly off of their respective non-target model using a smoothing factor of 0.01.
@@
 ADQN and DQN used a batch size of 32, whereas the RDQN, GRUDQN, and LSTMDQN models used a batch update by stacking a single trajectory.
-Other hyper parameters include $\gamma=0.95$ and $\epsilon=0.1$.
+Other hyper parameters included $\gamma=0.95$ and $\epsilon=0.1$.
 Machine specifications used in the experiment: 32 core Intel(R) Xeon(R) CPU E5-2640 v3 @ 2.60 GHz with 512 GB of Memory.
@@
 For all tasks in Experiment 3, we used $L:S\rightarrow 2^{P}$ and $j\geq i$ toggle construction.
-Hyper parameters include $\gamma=0.95$, $\alpha=0.1$ and $\epsilon=0.1$.
-We applied a similar sliding window as we did in Experiment 1 but under a lens of 50,000 episodes.
-We included an extended version of Table 1 in Figure 7.
+Hyper parameters included $\gamma=0.95$, $\alpha=0.1$ and $\epsilon=0.1$.
+We applied a similar sliding window, as we did in Experiment 1, but under a lens of 50,000 episodes.
+We include an extended version of Table 1 in Figure 7.
 Machine specifications used in the experiment: 32 core Intel(R) Xeon(R) CPU E5-2640 v3 @ 2.60 GHz with 512 GB of Memory.
-Figure 7: Results of ARMDPQ-Learning evaluated on Tasks (b) and (c) from the Breakfastworld domain on 10 random trials.
+Figure 7: Results of ARMDPQ-Learning evaluated over 10 random trials on Tasks (b) and (c) from the Breakfastworld domain.
 Task (b) and (c) Full show results when using the full RM for Task (b) and (c).
 Task (b) Cumulative shows results for using the RM for Task (b) where all rewards are 0 except the terminal cumulative rewards.
-Episode Reward Avg shows average reward +/- a stddev.
-$|U|$ shows current assumed size of $U$.
-$|T_{o}|$ shows current size of $T_{o}$.
-Cumulative Solve Time(s) shows the total solve time, for all ARMDPs, in seconds.
+Average Rewards column is displayed with standard deviations.
+$|U|$ column shows current assumed size of $U$.
+$|T_{o}|$ column shows current size of $T_{o}$.
+Cumulative Solve Time(s) represents the total solve time for all ARMDP inferences, in seconds.
+Appendix G Code
+All code available at https://github.com/GregHydeDartmouth/ARMDP_RL.git